### v2

# The Necessity of Imperfection: Reversing Model Collapse via Simulating Cognitive Boundedness

Abstract
Although synthetic data is widely promoted as a remedy, its prevailing production paradigm—one optimizing for statistical smoothness—systematically removes the long-tail, cognitively grounded irregularities that characterize human text. Prolonged training on such statistically optimal but cognitively impoverished data accelerates model collapse.
This paper proposes a paradigm shift: instead of imitating the surface properties of data, we simulate the cognitive processes that generate human text. We introduce the Prompt-driven Cognitive Computing Framework (PMCSF), whose core consists of a Cognitive State Decoder (CSD) that reverse-engineers unstructured text into structured cognitive vectors, and a Cognitive Text Encoder (CTE) that re-materializes these states into text enriched with human-typical imperfections via mathematically defined Cognitive Perturbation Operators.
The framework is validated through a two-stage objective evaluation pipeline. In cognitive codec verification, CTE text yields a Jensen–Shannon divergence of 0.0614 from human text (vs. 0.4431 for standard LLM output), passes double-blind professional media review, and achieves an intraclass correlation coefficient $ICC>0.9$ for cognitive profile alignment across heterogeneous models. In functional gain evaluation, isomorphic stress tests in the A-share market show that strategies incorporating CTE-generated data reduce maximum drawdown by 47.4% during the 2015 crash and deliver 8.6% Defensive Alpha, exceeding transaction costs by a factor of 33.
Our findings demonstrate that modelling human cognitive limitations—not copying surface data—enables synthetic data with genuine functional gain, offering a viable technical pathway toward resolving the AI data-collapse crisis.
Keywords: Synthetic Data; Model Collapse; Bounded Rationality; Cognitive Codec; Cognitive Latent Space Alignment; Cognitive Inter-lingua
1 Introduction
1.1 Structural Bottlenecks in AI Development: Statistical Convergence and the Loss of Cognitive Texture
Central to the evolution of artificial intelligence, and most notably within the domain of large language models (LLMs), lies a structural paradox; specifically, the consumption rate of anthropogenic data, upon which algorithmic advancement relies, vastly outstrips its natural genesis. Implicit in this disparity is the likelihood that high-quality human data reserves approach exhaustion [1]; concurrently, synthetic data, previously hypothesized as a viable substitute, has manifested as a primary catalyst for model collapse [3].
Situated at the root of model collapse are the fundamental distinctions between generative mechanisms; standard LLMs, relying principally on autoregressive probability prediction, gravitate toward statistically optimal solutions—outputs strictly aligned with the highest probability within the training distribution. This logic prioritizes fluency and convergence, thereby systematically excising traits inherent to human language, specifically those shaped by cognitive load, emotional oscillation, and heuristics and biases [14],[15].
It remains, however, precisely these deviations from optimality, termed herein as cognitive texture, that encapsulate the diversity and authenticity of human cognition. The utilization of denoised, AI-generated text for recursive training precipitates the erasure of long-tail cognitive information; consequently, a gradual homogenization of behavioral patterns ensues. Adopted within this research is a human-AI collaborative R&D paradigm designed to expedite a 0-to-1 proof-of-concept under resource constraints; such an approach delineates a potential trajectory for the field.
Supplementary materials, including raw data logs and deployed parameter ranges, are available in the ancillary files of this arXiv submission. Please refer to these files or contact the author for further details.
1.2 The Core Breakthrough: Simulating the “Cause” Behind the Cognitive Mirror
To mitigate the aforementioned dilemma, this study proposes a paradigm shift; effective synthetic data generation, rather than ceasing at the statistical imitation of surface appearance, must transition to a deep simulation of the cause underlying real-world text, namely, the bounded rationality governing human decision-making [4], [5].
Foundational to this inquiry is the Cognitive Mirror Hypothesis; positing that, as human text emerges from constrained cognitive resources, the cognitive state responsible for its generation remains theoretically reversible. Should the reverse-engineering, or decoding, of structured cognitive states from text prove feasible, it appears plausible that forward engineering, or encoding, could reconstruct text from these states under controlled cognitive constraints.
Predicated on this premise, the requirement for AI to pursue statistically optimal outputs becomes obsolete. Delineated instead is a constraining framework that compels the simulation of a human agent attempting to produce a satisficing, or merely adequate, response under cognitively limited conditions; included within this simulation are the non-linearities influenced by biological neural noise.
1.3 Methodology Preview: The PMCSF Framework and Two-Stage Objective Validation
Building on this foundation, the present study constructs a Prompt-driven Cognitive Computing Framework (PMCSF)—a dual-engine system engineered for inverse operations that encompasses two core components:
  • Cognitive State Decoder (CSD): Tasked with the inverse mapping of cognitive states, this component translates unstructured natural language text into a 17-dimensional cognitive state vector space via dimensionality reduction, thereby achieving a structured representation of latent psychological traits.
  • Cognitive Text Encoder (CTE): Charged with the forward simulation of bounded rationality, this component leverages a dual-layer cognitive simulation architecture: the macro layer internalizes cognitive priors to construct decision-making scenarios constrained by limited mental capacity or time pressure, while the micro layer invokes cognitive perturbation operators to systematically introduce irrational fluctuations.
To validate the efficacy of this paradigm, the study devises a progressive validation strategy spanning morphological fidelity to functional gain:
  • Cognitive Codec Mechanism Validation: Employs multi-dimensional validation via double-blind ecological tests, computational linguistic analysis, and cross-model consistency tests (N=26) to assess the decoding consistency of the CSD and the encoding fidelity of the CTE.
  • Functional Gain Validation: Subjects the synthetic data to stress testing within the A-share market, a complex system characterized by high-frequency feedback; given the financial market’s extreme sensitivity to emotional noise and irrational behavior, this test aims to verify whether the synthetic data effectively complements the human long-tail cognitive signals filtered out by mainstream models—a hypothesis for which the financial market’s sensitivity to emotional noise and irrationality serves as an optimal crucible.
1.4 Contributions of This Study
The primary contribution of this research resides in challenging the mainstream rational paradigm of AI development, proposing and empirically validating a new design principle: “Controlled Non-Optimality.” Contributions span three primary levels:
  • Theoretical Innovation: The PMCSF posits a 17-dimensional affective vector space as an intermediate cognitive language shared by humans and machines, elucidating that irrational thinking is not incalculable noise but high-dimensional information encapsulating “Cognitive Invariants”—structures amenable to representation and reconstruction.
  • Methodological Innovation: Constructs a semantic projection protocol that decodes the implicit, high-dimensional machine cognition within models into an explicit 17-dimensional human cognitive coordinate system, thereby forging a pathway between LLMs and human thought/emotion.
  • Empirical Innovation: Furnishes evidence for the “enabling value” of synthetic data within a high-fidelity simulation of the A-share market; under realistic constraints—including transaction costs and risk metrics—it demonstrates the risk management advantages and robustness of cognitive-enhanced synthetic data in extreme market environments.
2 Theoretical Foundations and Related Work
2.1 Starting Point: Data Exhaustion and “Statistical Mode Collapse”
Addressing the increasingly severe Curse of Recursion [3] plaguing contemporary AI development, this study takes root in a growing crisis of data reliability: as high-quality human data depletes [1], reliance on AI-generated synthetic data for model training has transitioned from a discretionary practice to an existential necessity. Yet the prevailing paradigm for synthetic data generation harbors a critical flaw: standard large language models (LLMs)—trained via Maximum Likelihood Estimation (MLE)—exhibit a systematic bias toward generating samples aligned with the statistical mode of their training distribution. This yields a predilection for text that is syntactically smoothed, statistically common, and grammatically conventional—traits that filter out the idiosyncrasies (e.g., hesitation, logical leaps) defining authentic human expression.
Concurrent with this, such formally standardized text is polluting contemporary content production. A 2024 NewsGuard report [2] documents thousands of websites leveraging AI to mass-produce low-quality content rife with factual errors; training models on such data, the study argues, will inevitably degrade the performance of next-generation systems [3, 18]. The root cause lies in LLMs’ pursuit of statistical optimality: by converging toward the mean of the training distribution, these models systematically erase the “noise”—the irregularities and cognitive texture—that distinguishes human language from algorithmic output. Over time, this process of Statistical Mode Collapse erodes the diversity of synthetic data, setting in motion the degenerative cycle of Model Collapse where AI systems lose the ability to represent the tails of the original data distribution.
2.2 Theoretical Cornerstone: Bounded Rationality as the Authentic Texture of Cognition
The detectability of AI-generated text by classification models stems not from inherent low quality but from an overweening structural perfection—a trait absent in authentic human expression. Human language, the study posits, derives its authenticity from Cognitive Non-Optimality—the imperfections arising from bounded cognitive resources, emotional flux, and heuristic reasoning. This insight rests on three foundational axioms from cognitive science:
  • Bounded Rationality [4, 13]: Nobel laureate Herbert Simon’s theory posits that humans, constrained by incomplete information, limited cognitive capacity, and temporal pressures, abandon the pursuit of optimal decision-making in favor of a satisficing strategy—accepting solutions that meet a threshold of adequacy rather than perfection. This “good enough” heuristic, the study argues, is not a flaw but a defining feature of human cognition.
  • Heuristics and Biases [5]: Research by Kahneman and Tversky demonstrates that humans rely on mental shortcuts (heuristics) for rapid judgment, yet these shortcuts generate systematic biases (e.g., loss aversion, recency bias). These biases, the study contends, introduce emotional fluctuations and perspectival narrowness into text—traits that standard AI systems filter out in their quest for statistical fluency.
  • Competence vs. Performance [6]: Noam Chomsky’s distinction between competence (idealized grammatical knowledge) and performance (error-prone, hesitant linguistic output) underpins a key critique of AI: while LLMs excel at replicating the rigid perfection of competence, they fail to capture the living, dynamic performance of human language.
Implicit in these findings is a paradox: bounded rationality and cognitive imperfections should not be treated as mere noise; they constitute the authentic texture of human cognition. For synthetic data to possess training value, the study argues, it must faithfully simulate these non-optimal characteristics—reproducing the “irrational” fluctuations that make human language both diverse and meaningful.
2.3 Implementation Pathway: LLMs as a Cognitive Simulation Platform
Within academic circles, a consensus is coalescing: LLMs operate not merely as text generators but as a universal cognitive simulation platform [16, 17]. The feasibility of this framing rests on two pillars:
  • Implicit Pattern Representation: Through massive pre-training on human text, the internal parameter space $\Theta_{i}$ of LLMs has formed a high-dimensional representation of human expression patterns—including the statistical features arising from bounded rationality. Yet under default generation strategies (e.g., top-k sampling), these features are systematically smoothed, rendering them invisible in standard outputs.
  • Cognitive Plasticity: By introducing external constraints—specifically, the PMCSF framework proposed in this study—researchers can guide the generation process to deviate from the path of statistical optimality. This activation of the latent distribution’s “irrational, highly emotional, low-probability” regions enables targeted simulation of specific cognitive states, transforming LLMs from mere text generators into tools for modeling human thought.
2.4 Operationalization Mechanism: Mathematical Compilation from Psychological Principles to Cognitive Perturbation Operators
Building on this theoretical framework, the study transmutes prompt engineering—long an empirical exercise in parameter tuning—into a rigorous methodology of Cognitive Mathematical Compilation. Each Cognitive Perturbation Operator in the PMCSF framework is a mathematical mapping of a specific cognitive psychology principle, designed to inject “biological noise” into the text generation process:
  • Sentence Length Oscillation Operator $L_{s}(n)$:
    $L_{s}(n)\propto\sin(\omega n)+\mu$ (1)
    Drawing on Baddeley’s Working Memory Model [8] and Sweller’s Cognitive Load Theory [9], this operator simulates the “load-release” cycles of human cognition by forcing sentence length to oscillate rhythmically between long, complex structures (high cognitive load) and short, declarative phrases (low load). The sinusoidal function, the study argues, mirrors the biological constraints of breathing and attention, introducing a “natural” variability absent in standard AI output.
  • Probability Perturbation Operator $f_{w}(t)$:
    $f_{w}(t)\approx p(t)\times(1-\beta)+\epsilon$ (2)
    To simulate human “hesitation” and “non-optimal choice,” this operator introduces controlled noise $\epsilon$ into the token selection process, forcing the model to deviate from the probability peaks of the training distribution. The result is text that retains the viability of long-tail vocabulary—words and phrases that, while statistically unlikely, are central to authentic human expression.
  • Associative Leap Operator:
    $\cos(\theta_{context},\theta_{word})<0.5$ (3)
    Corresponding to nonlinear associative leaps in human thought (e.g., topic shifts triggered by external stimuli), this operator sets a semantic distance threshold for token generation. By allowing tokens with low cosine similarity to the current context, the model breaks free from its inherent linear logical inertia, producing text that mirrors the “unpredictable” jumps of human dialogue.
Collectively, these operators form the micro-foundation of the CTE (Cognitive Text Encoder)—a component designed to ensure synthetic data is not random gibberish but text that conforms to biological cognitive patterns. The study grounds this design in Gigerenzer’s theory of adaptive toolboxes [24], which posits that human decision-making in complex environments relies on “fast and frugal” heuristics rather than complex Bayesian optimization. The perturbation operators, therefore, are not mere noise; they are mathematical simulations of the Ecological Rationality inherent in human thinking—adaptive imperfections that enable survival in uncertain environments.
3 The Prompt-driven Cognitive Computing Framework
3.1 Framework Design Philosophy: A Human-AI Universal “Cognitive Codec Protocol”
Predicated on a generative architecture which draws its theoretical validity from the Cognitive Mirror Hypothesis, this study explicates the Prompt-driven Cognitive Computing Framework (PMCSF). Designed to mitigate the erosion of Cognitive Texture in standard LLMs—an artifact of their systematic prioritization of statistical smoothing—the framework establishes a bidirectional mapping protocol between natural language and cognitive states. Central to this initiative is the computable representation of human Bounded Rationality, a core objective underpinned by the hypothesis that cognitive states and linguistic outputs possess reversible mappability.
Consequently, PMCSF functions as a universal Human-AI Cognitive Codec protocol, comprising two inverse operational engines:
  • Cognitive State Decoder (CSD): Functioning as the “reading” engine (inverse engineering), this component maps unstructured human text onto a 17-dimensional cognitive state vector space via dimensionality reduction. This process, which extracts the underlying cognitive state, translates the ambiguity of natural language into a structured, machine-interpretable format.
  • Cognitive Text Encoder (CTE): Serving as the “writing” engine (forward engineering), the CTE reconstructs structured cognitive states into text imbued with human characteristics under the constraints of controlled bounded rationality. Anchored by a dual-layer architecture—macro-anchoring for satisficing and micro-perturbation for cognitive noise—this design distinguishes PMCSF from conventional generative models, which typically prioritize statistical optimality over authentic linguistic variation.
Figure 1: Architecture of the Prompt-driven Cognitive Computing Framework (PMCSF). Operating as a “Cognitive Codec”, the system incorporates two inverse engines: CSD (dimensionality reduction) and CTE (reconstruction). The diagram delineates the invariant 17-Dim Cognitive State Vector serving as the intermediate language, alongside CTE’s dual-layer design (Macro-Anchoring & Micro-Perturbation) for simulating bounded rationality. This topology—grounded in the hypothesis of cognitive invariants—suggests structural stability across diverse AI models and contexts.
3.2 Component 1: Cognitive State Decoder (CSD): The Inverse Engineering of Cognition
Addressing the challenge of interpreting human linguistic inputs through the quantification of ambiguous natural language, the CSD diverges from standard LLMs. While conventional models prioritize statistical likelihood over cognitive fidelity, the CSD is engineered to capture the “tails” of linguistic variation—those idiosyncratic expressions that reflect bounded rationality and emotional flux.
3.2.1 Construction of the 17-Dimensional Cognitive State Vector Space
Synthesizing discrete and dimensional emotion models, this study developed a 17-dimensional vector space grounded in cognitive psychology axioms, which serves as the system’s Cognitive Inter-lingua. Structured to capture both universal and domain-specific states, the dimensions encompass:
  • Basic Emotion Layer: Drawing on Ekman’s discrete emotion model [21], this layer incorporates core basis vectors including joy, sadness, anger, fear, trust, disgust, surprise, and anticipation—states posited to be invariant across cultural and linguistic contexts.
  • Cognitive Regulation Layer: Leveraging Russell’s dimensional emotion model [22], this layer integrates constructs such as intensity (Arousal), agency (Dominance), certainty, and temporality, which modulate how basic emotions are manifested in language.
  • Domain-Specialized Layer: Tailored to the financial trading context, this layer includes higher-order cognitive states such as FOMO (Fear Of Missing Out), Greed, Regret, and Uncertainty—states that drive market behavior yet are frequently filtered out by standard LLMs due to their statistical rarity.
Figure 2: 17-Dimensional Cognitive State Vector Comparison. Radar chart visualizing distinct “Cognitive Topologies” decoded by CSD. Note the structural divergence between Panic Mode (Red; high Fear/Uncertainty) and Frenzy Mode (Blue; high Greed/FOMO). This confirms CSD’s capability to disentangle complex market sentiments into interpretable mathematical vectors.
3.2.2 Quantification Mechanism: Constrained Prompt-based Probabilistic Projection
Harnessing the implicit statistical representation of natural language patterns encoded by LLMs, this study reframes latent syntactic and semantic structures into a tool for measuring probabilistic distributions via prompt engineering. Within this process, the system imposes targeted constraints that compel the model to project unstructured text onto the 17 orthogonal cognitive dimensions. Generating corresponding confidence scores, this mechanism enables the high-precision decoupling and quantification of latent cognitive states embedded within human language. (See Appendix B.1 for the complete constrained prompt pseudocode.)
Drawing on empirical findings, this study substantiates that purely descriptive natural language prompts—which lack structured cognitive constraints—exhibit significant "semantic drift" across divergent models and temporal frames. Mathematical operators (e.g., Lₛ(n)) are introduced not to prioritize mechanical precision, but to act as high-dimensional cognitive anchors that lock the structural skeleton of text generation in latent space, thereby ensuring high cross-model consistency.
3.2.3 Macro-Aggregation & Dynamic Prediction
To account for the non-linear dynamics inherent in market behavior, the CSD extends its functionality by constructing a multi-level cognitive processing flow:
  • Macro State Assessment (Node 2): Consolidating daily micro cognitive vectors into a macro state vector ($V_{state}=\{MDI,MCFI,Meta\}$), this node calculates the Market Dispersion Index (MDI) to quantify sentiment divergence and the Market Consensus Frenzy Index (MCFI) to measure irrational exuberance. These metrics serve as the foundational basis for identifying market regime shifts.
  • Context-Adaptive Prediction (Node 3): Drawing on the macro quadrant probabilities output by Node 2, the framework employs a hybrid-precision GJR-GARCH architecture. Within this paradigm, the model eschews fixed parameters, instead dynamically modulating volatility transmission coefficients ($\alpha$, $\beta$) in real-time response to the current market state (e.g., “Structural Tearing”). This adaptive modulation facilitates the capture of non-linear risk signals that typically elude standard static models.
3.3 Dual-layer Cognitive Simulation Framework
Rooted in a fundamental redefinition of the generation objective, the CTE (Cognitive Text Encoder) shifts its optimization focus away from textual fluency toward the systematic reproduction of human cognitive imperfections. Designed to circumvent the statistical smoothing inherent to autoregressive models, the study advances a dual-layer cognitive architecture:
  • Macro Layer: Aligned with Simon’s Satisficing paradigm, this layer internalizes “Cognitive Priors” to simulate the “Bounded Search” process characteristic of human decision-making. Constrained by knowledge boundaries and perspectival limits, this mechanism forestalls the convergence of generated content toward the statistical mean, thereby preserving the specificity of individual viewpoints.
  • Micro Layer: Corresponding to Baddeley’s working memory construct and Kahneman’s noise hypothesis, this layer deploys Cognitive Perturbation Operators. Including functions such as the “Sentence Length Oscillation Operator” and the “Probability Perturbation Operator,” this layer injects systematic irrational fluctuations at the lexical and syntactic levels. Furthermore, a “Logical Fault Tolerance Mechanism” permits micro-level logical leaps, reproducing the instability of biological neural networks [10].
At the level of underlying operational mechanisms, the dual-layer architecture of the PMCSF replicates the core mechanism of human creativity: conceptual integration [23]. Whereas advanced human creation entails the decoupling and recombination of discrete cognitive elements to generate novel syntheses, standard AI tends to smooth these heterogeneous elements. By orthogonally separating top-level macro anchoring from bottom-level micro perturbation, the PMCSF enables the systematic recombination of these elements, fostering the emergence of cognitive tension.
3.4 Core Principles of the Framework: Anti-Data Leakage and Causal Simulation
Rigorous adherence to the Anti-Data Leakage Principle governs the design of the PMCSF. During construction, the system strips away all specific historical facts and objective market summaries that could induce bias, retaining only universal commentary information. Combining this stripped-down dataset with relevant mathematical theories, the system establishes cognitive dynamics rules. These principles ensure the framework operates as a general cognitive dynamics simulator rather than an overfitted repository of historical data.
4 Experimental Design and Data
4.1 Philosophy of Experimental Design: From Statistical Fitting to Causal Prototype Verification
Rigorously validating the efficacy of the PMCSF framework, this study constructs a progressive two-stage verification system. While accounting for financial market non-stationarity—a challenge highlighted by Harvey [19]’s warnings about “p-value manipulation” and López de Prado [20]’s framework for the “non-stationarity dilemma”—the experimental design abandons the traditional quantitative research approach of “large-sample statistical fitting” in favor of a prototype verification strategy prioritizing “causality over correlation.” This shift reflects a deliberate rejection of correlation-driven inference, which risks conflating spurious patterns with actionable insights in dynamic systems.
4.1.1 Sample Selection Under Non-Stationarity Constraints
Training on mixed-period data, as both scholars caution, appears prone to introducing structural noise that obfuscates causal relationships. Accordingly, this study eschews full historical data in favor of a “Prototype Verification” principle, selecting “Scenario Anchors” from A-share market history—specifically the 2015 stock market crash and the 2024 bull market—which exhibit the highest signal-to-noise ratios and most pronounced irrational characteristics as experimental samples. Employing high-precision sample validation, this approach aims to confirm the model’s robustness and logical self-consistency under extreme market conditions, where cognitive biases and information asymmetry are most acute.
4.1.2 Phased Verification Process
Guided by the foregoing principles, the verification process is partitioned into two stages, each designed to isolate distinct dimensions of the PMCSF framework’s performance:
Phase One: Verification of Cognitive Codec Mechanisms
It aims to substantiate the objectivity of PMCSF as a universal cognitive protocol—an assertion contingent on its ability to encode and decode human language consistently across contexts. To this end, a three-dimensional testing framework, which includes double-blind testing (to eliminate observer bias), Jensen–Shannon divergence (to quantify distributional similarity), and cross-model testing (to assess generalizability), is constructed to comprehensively validate the operational efficacy of the dual-engine architecture (CSD/CTE).
Phase Two: Functional Gain Verification
Under live simulated stress testing in the A-share market—a setup mimicking the constraints of real-world trading—the phase assesses whether CTE-generated data effectively complements the long-tail cognitive signals filtered out by standard AI. These signals, which include rare heuristic biases and emotional fluctuations, are critical to capturing the “cognitive texture” of human decision-making. The goal is to determine whether such data can yield substantial strategic improvement, particularly in environments where statistical optimality (e.g., mode collapse) undermines predictive accuracy.
4.2 Phase I: Verification of the Cognitive Codec Mechanism
Tasked with validating the foundational mechanisms of the PMCSF framework, this stage aims to comprehensively substantiate that the system not only generates text imbued with human-like cognitive texture but also captures objective cognitive invariants—core constructs posited to underpin authentic human language.
4.2.1 Principle Verification: Double-Blind Ecological Test and Algorithmic Feedback
In pursuit of ecological validity, in-situ tests were conducted over a one-month period (August–September 2025) within two authoritative Chinese internet content distribution environments. Guided by academic ethics and privacy protections, the platforms are anonymized:
Platform A (Expert Human Review Environment)
  • Characteristics: One of China’s most influential top-tier technology and business media platforms, which maintains stringent entry barriers (average external submission acceptance rate $10\%$ over the past three years) and requires all external submissions to undergo a triple-review process by senior editors (with >5>5 years of experience).
  • Experimental Design: A double-blind design was employed, wherein submissions from $D_{CTE}$ (generated by this framework) and $D_{Human}$ (authored by human experts) were mixed and submitted over one month. Review editors, completely unaware of the experiment, selected and published articles based on their professional experience.
  • Evaluation Metric: Submission acceptance rate, a metric that serves as the highest standard of human expert aesthetic judgment.
Platform B (Algorithm-Human Hybrid Recommendation Environment)
  • Characteristics: A comprehensive news and information portal boasting a stable monthly active user (MAU) base exceeding 200 million.
  • Distribution Mechanism: The portal employs a hybrid distribution model combining deep learning recommendation algorithms and human intervention, where content first undergoes algorithmic initial quality screening to enter a basic traffic pool before receiving traffic weighting based on user interaction metrics (e.g., click-through rate [CTR], read completion rate).
  • Evaluation Metric: Average views per article, a measure that reflects objective feedback on content quality from an industrial-grade recommendation algorithm.
4.2.2 Statistical Fingerprint Verification: JS Divergence Metrics
In an effort to eliminate subjective evaluation bias, computational linguistics metrics were introduced to objectively verify whether CTE-generated text ($D_{CTE}$) exhibits a statistical distribution closer to real human text ($D_{Human}$) than does standard AI-generated text ($D_{Standard}$).
  • Design:
    • Sample Construction: To ensure comparability, three parallel corpora were constructed for the same market period (e.g., the 2015 stock market crash), including (1) Real human comments ($D_{Human},N=100$); (2) CTE-generated text ($D_{CTE},N=100$); (3) Standard AI-generated text ($D_{Standard},N=100$).
    • Metric Selection: Drawing on Biber’s multi-dimensional analysis framework [7], key statistical features—including sentence length standard deviation and adjective density—were selected as linguistic fingerprints, with commas used to denote categorical boundaries.
  • Evaluation: Jensen-Shannon (JS) Divergence, a measure of distributional similarity, was used to quantify the distance between the probability distributions of each generated group and the human benchmark ($D_{Human}$).
  • Inference: If $JS(D_{CTE}||D_{Human})\ll JS(D_{Standard}||D_{Human})$, the data lends credence to the claim that the CTE successfully breaks the statistical smoothing of standard AI and replicates the cognitive texture of human language at a structural level.
  • Note: $D_{Standard}$ was generated using the advanced “Immersive Market Participant Simulator” protocol (see Appendix H), which enforces specific persona constraints (e.g., “bagholder,” “hot money”) and expression rules to minimize AI-typical neutrality, representing a rigorous baseline for state-of-the-art prompt engineering.
4.2.3 Biological Volatility Verification
Seeking to transcend the constraints of subjective Turing tests and probe the latent mechanisms governing model-generated text, this study adopts a micro-statistical lens—conducting normality tests and dispersion analyses on the sentence-length distributions of synthetic outputs. Central to this section is the objective of employing quantitative methodologies to rigorously validate the empirical validity of the “Sentence Length Oscillation Operator $L_{s}(n)$” at the micro-statistical scale.
  • Design:
    • Sample Construction:
        Control Group (Standard-AI): A cohort of 155 sentences ($N=155$) was extracted from output generated by an industry-recognized high-quality model (Standard-AI). Declaration: This sample underwent independent assessment by a senior editorial team affiliated with a leading authoritative Chinese financial media outlet—an evaluation that substantiated compliance with publishing standards for grammatical accuracy, logical coherence, and terminological precision, with the text deemed “indistinguishable from the work of a human senior commentator.” This rigorous endorsement ensures the study avoids comparison with suboptimal models, instead targeting the industrial “state-of-the-art” in text generation.
        Experimental Groups (CTE Groups): Drawn from a text pool generated through double-blind testing, three representative texts—randomly selected to represent divergent scenarios—were designated as CTE-A (Emotional Catharsis Mode), CTE-B (In-depth Analysis Mode), and CTE-C (Macro-Chaotic Mode).
        Sample Construction Strategy: Employing a sentence-level granularity analysis strategy, this study conducts comparisons across four representative text scenarios while deriving statistical power from the micro-units within these texts. The total sample comprises 461 sentences (Standard Group: 155; CTE Groups: 113/90/103 respectively), a dataset that surpasses the statistical threshold required for robust distribution analysis and Shapiro-Wilk normality tests (N >> 30). The observed Zipfian long-tail distribution and high variance emerge as structurally significant features with statistical robustness, a conclusion supported by the dataset’s adherence to the thresholds for normality and distributional integrity.
    • Data Processing: Full texts from both the Standard-AI and CTE groups were parsed, with segmentation into granular units performed using standard Chinese sentence-ending punctuation (period, question mark, exclamation mark); subsequent analyses quantified character length for each sentence and total sentence count.
    • Metric Selection: To capture central tendency, the Mean ($\mu$) was selected; for relative dispersion—dubbed “breathing sensation”—the Coefficient of Variation (CV) was employed. Distribution shape, particularly long-tail characteristics, was assessed via Skewness, while normality was evaluated using the Shapiro-Wilk Test.
  • Evaluation: The distribution of Standard-AI outputs ($N=155$) served as the benchmark, with its probability density function computed to assess conformity to normality ($p>0.05$) and coefficient of variation. A distribution meeting the normality assumption while displaying low dispersion was classified as having succumbed to the “Trap of Statistical Smoothness.” Statistical characteristics of the CTE groups were then compared to this benchmark. Should the CTE data exhibit significantly elevated CV, right skewness (indicating long-tail distributions), and a rejection of normality ($p<0.05$), such outcomes would be interpreted as successful activation of biological volatility.
  • Inference: If experimental results reveal Standard-AI outputs converging to a normal distribution, whereas CTE groups display non-normal, high-volatility characteristics with statistically significant differences, the data would lend credence to the hypothesis that the CTE effectively disrupts the “statistical smoothness” inherent to standard AI—replicating, at the structural level, the “quasi-periodic chaos” and Cognitive Texture that define authentic human language.
4.2.4 Verification of Cognitive Structure Objectivity
Objective: To verify whether the proposed 17-dimensional cognitive state vector space—a construct designed to represent human cognitive processes—possesses objective scientific validity independent of model architecture.
Logic: Positing that a subjectively defined vector space would produce random deviations in quantitative data across models, the study argues that high structural convergence among heterogeneous models (with different architectures, e.g., DeepSeek V3 vs. Doubao 1.6) within this framework would confirm the capture of objective cognitive topology and cognitive invariants inherent in human language.
Design: A test set containing 26 key time nodes was constructed, and the Intraclass Correlation Coefficient (ICC) for hierarchical sentiment indices between the two models was calculated to quantify their consensus on the cognitive structure.
4.3 Phase II: Functional Gain Verification (A-Share Market Hypothesis)
Building upon the morphological fidelity verification, this phase deploys the PMCSF framework within the A-share market ecosystem, aiming to evaluate the signal completeness efficacy of CTE data in complex financial systems via a quantitative backtesting paradigm. The experimental hypothesis posits that Cognitive-Enhanced Data—by injecting irrational cognitive factors—can refine a model’s analytical prowess regarding market microstructure, a claim rooted in the premise that human-like “imperfections” strengthen rather than obscure decision-making signals.
To mitigate theoretical backtesting biases, the study incorporates stringent constraints congruent with live trading standards:
  • Transaction Friction: A total cost of 0.26% per trade—encompassing bilateral commissions and impact slippage—was established to replicate real-world execution costs.
  • Risk Benchmark: An annualized risk-free rate of 2% (daily 0.008%) was designated for Sharpe Ratio calculation, aligning with industry-standard risk metrics.
  • Evaluation System: Primary focus was directed toward Maximum Drawdown and the Sharpe Ratio (beyond cumulative returns) to assess strategy robustness under extreme risk conditions.
4.3.1 CSD Interpretability Verification: Four-Stage High-Fidelity Protocol
To address the methodological challenge of inconsistent time granularity in the prototype sample, the study adopted a four-stage hierarchical verification protocol:
  1. 1.
    Stage 1 (Quantitative Macro Signals): For a sample of $N=16$ “event days,” the Pearson correlation between the CSD sentiment index and the objective index’s daily price change was tested to validate directional alignment.
  2. 2.
    Stage 2 (Qualitative Narrative Verification): Across the full sample of $N=27$, a frame-by-frame comparison was conducted between the core narrative extracted by the CSD and the semantic matching degree with objective event descriptions, ensuring narrative coherence.
  3. 3.
    Stage 3 (Causal Mechanism Verification): Key prototypes were spot-checked to compare whether the cognitive biases and causal chains diagnosed by the CSD aligned with objective descriptions of market behavior.
  4. 4.
    Stage 4 (Core Feature Verification): The CSD’s most distinct function—investor stratification—was verified to determine if its results quantitatively replicated real historical complex divergences (e.g., “retail investor bear market vs. fund investor bull market”), a test of cross-temporal consistency.
4.3.2 Dynamic Prediction Superiority Verification: M-Dynamic vs. M-Static
Objective: To ascertain whether the CSD-based “context-adaptive” model (M-Dynamic) outperforms a static benchmark with fixed parameters (M-Static), a comparison designed to isolate the value of real-time cognitive adaptation.
Design: Comparative tests were conducted during two independent out-of-sample (OOS) periods—April 2025 and November 2025—to avoid overfitting to a single market regime.
  • Experimental Group (M-Dynamic): Employed the full context-adaptive model, leveraging CSD-derived signals to adjust parameters in response to market flux.
  • Control Group (M-Static): Deployed a static benchmark model with parameters fixed to average values across all quadrants, simulating a “frozen” decision-making framework.
  • Evaluation Dimensions: Encompassed both the mechanistic layer (signal response speed, information entropy) and the practical layer (Safety Buffer, Sharpe Ratio) to capture trade-offs between theoretical rigor and real-world utility.
4.3.3 Empowerment Value Verification: The “Trace Element” Effect of CTE Data
Aiming to verify the core value of CTE data as a “cognitive enhancer,” the experiment constructs three comparison groups to test the effectiveness (Information Coefficient, IC) of the sentiment factor during two extreme cycles: the 2015 stock market crash ($N=23$) and the 2024 bull market ($N=13$):
  • Model A (Cognitive-Enhanced): Incorporated 20% CTE data, blending human and synthetic signals to simulate the “noise trader” dynamics described in the DSSW Model.
  • Model B (Pure Human): Relied on 100% pure human data, serving as a baseline for assessing signal dilution risks.
  • Model C (Standard AI): Integrated 20% standard AI data, a control to measure the impact of statistically optimized (but cognitively flat) inputs.
Note: Drawing on the DSSW Model (De Long et al.) from behavioral finance, the 20% blending ratio was chosen to mimic the mechanism whereby a minority of irrational traders dominate peripheral pricing. This design mirrors the bi-domain integration in conceptual blending theory [23], seeking to enhance the model’s generalized understanding of non-stationary financial signals by injecting heterogeneous Cognitive Texture.
Inference: If $IC_{A}>IC_{B}>IC_{C}$, the data would lend credence to the claim that CTE data not only avoids diluting the human signal but—through the complementation of irrational cognitive features—enhances the model’s capability to capture extreme market conditions, a finding with direct implications for risk management.
4.3.4 Isomorphic Stress Testing Based on Real-World Simulation
To validate that CTE data effectively captures long-tail cognitive signals overlooked by standard data, the study designed a rigorous isomorphic ablation experiment—one that holds all variables constant except for the presence of cognitive decision triggers:
  • Experimental Logic and Variable Control: The base strategy was uniformly set as the M-Dynamic framework with GARCH parameter adaptability, designating the intervention state of the “Conditional Trigger Logic” as the sole controlled variable.
  • Experimental Group (20% CTE Data): Introduced an explicit threshold mechanism based on the CTE cognitive state vector (e.g., setting $fear>0.3$ as a mandatory stop-loss point), a setup intended to simulate a trading execution process with clear cognitive decisiveness.
  • Control Group (100% Human Data): Removed the cognitive threshold trigger module, relying solely on statistical mean reversion logic to simulate a passive adaptation state devoid of a clear action trigger.
Verification Goal: The experiment seeks to determine whether the high signal-to-noise ratio information from CTE data can convert a potential information perception advantage into precise timing capability and strategic gains at key market inflection points (e.g., trend reversals, structural breakdowns) during non-linear market conditions like the 2015 crash and 2024 bull market.
5 Experimental Results
Systematically presenting empirical data from the PMCSF framework across two dimensions—cognitive encoding/decoding mechanisms (per the “Cognitive Codec” definition) and functional gain—this chapter demonstrates that the framework not only approximates the statistical fingerprints of human cognitive texture (the “imperfections” inherent in authentic human language) but also yields significant asymmetric strategic advantages in high-frequency feedback financial systems.
5.1 Phase I: Verification Results of the Cognitive Codec Mechanism
This phase validates the fundamental operational capabilities of the dual-engine architecture comprising the CTE (Cognitive Text Encoder) and CSD (Cognitive State Decoder)—the core engines of the PMCSF framework.
5.1.1 Industrial-Grade Turing Test: Sensory Indistinguishability
Employing a double-blind experimental design, this study confirmed that the PMCSF framework generates long-form text indistinguishable from human-written content at a sensory level—findings that lend preliminary validation to the Simulating Cognitive Non-Optimality principle.
Expert Review Pass Rate (Platform A)
Over a one-month period, during which reviewing editors remained blind to content provenance, CTE-generated submissions achieved a 72.7% expert review pass rate—significantly outperforming the 13.1% rate of the human control group ($p<0.01$). Given that this platform typically rejects over 80% of human submissions, the results suggest that cognitively perturbed text—defined as writing bearing the “imperfections” of cognitive load and emotional fluctuation—did not merely meet but exceeded the baseline level of human experts in logical density and professional insight.
Algorithmic Recommendation Feedback (Platform B)
Under controlled conditions ensuring topic consistency, CTE-generated data outperformed the human group, with the average read count per CTE-generated article (11,089) significantly exceeding that of the human group (7,314). This outcome is attributed to the cognitive non-optimality simulated by the CTE—including emotional fluctuations, heuristic biases, hesitation, logical leaps, and controlled text rhythm—which effectively triggered positive feedback mechanisms in the recommendation algorithm. The content was judged to possess higher user retention value, resulting in greater traffic weighting.
5.1.2 Morphological Fidelity: Statistical Fingerprint Verification Based on JS Divergence
Seeking to mitigate subjective bias, this study employs computational linguistic metrics—rooted in Biber’s [7] multidimensional analysis framework—and deploys Jensen-Shannon (JS) divergence to quantify the distributional distance between $D_{CTE}$ (CTE-generated data) and $D_{Human}$ (human data) across five key statistical dimensions. As delineated in Table 1, CTE-generated data exhibit a pronounced statistical overlap with human data when evaluated against stylistic fingerprint metrics.
  • Key Finding: For the standard deviation of sentence length, the Jensen-Shannon divergence between $D_{CTE}$ and $D_{Human}$ registers at a mere 0.0614, while the corresponding value for standard AI-generated data ($D_{Standard}$) reaches 0.4431.
  • Theoretical Interpretation: This marked (>7>7-fold) discrepancy lends credence to the efficacy of cognitive perturbation operators. Consider, for example, the sentence length oscillation operator $L_{s}(n)$: it reliably disrupts the statistical smoothing (i.e., Statistical Mode Collapse) characteristic of standard large language models, thereby replicating the cognitive “breathing” pattern inherent to human linguistic production.
Table 1: Comparison of JS Divergence in Statistical Style Fingerprints
Metric CTE vs Human
(Std-AI vs Human) Key Insight
Sentence Length SD 0.0614
0.4431 CTE approximates human (>7>7x diff).
Adjective Density 0.0526
0.6094 CTE approximates human (>11>11x diff).
Noun-Verb Ratio 0.1197
0.2882 CTE is significantly closer to human.
Interjection Count 0.0003
0.0101 CTE is significantly closer to human.
Avg Sentence Length 0.0935
0.5054 CTE approximates human (>5>5x diff).
Note: Top row is CTE data; Bottom row (italic) is Std-AI data.
5.1.3 The Statistical Smoothness Trap and Biological Fluctuation
Adopting a micro-statistical lens, this study conducts normality tests and dispersion analyses on the sentence length distributions of generated text, with visualizations presented in Figure 3.
Figure 3: The Spectrum of Cognitive Rhythm: Statistical Smoothness vs. Biological Fluctuation. (A) Distribution: Standard AI (blue) remains confined to a symmetric normal distribution, while CTE models (red/green/purple) demonstrate human-like right-skewness and Zipfian tails. (B) Variability: Box plots reveal the rigid stability of Standard AI ($CV\approx 44\%$) alongside the high dynamic range and outliers of CTE models ($CV\approx 65\%$), thereby verifying the Sentence Length Oscillation Operator. (C) Fingerprints: The radar chart contrasts the static nature of Standard AI (centered) with the dynamic, context-dependent signatures of CTE models (expanded).
The Trap of Statistical Smoothness
As depicted in Figure 3, the sentence length distribution of Standard AI (Group 1) manifests an unnatural “perfect symmetry.”
  • Data Characteristics: Its data, clustering tightly around a mean of 23.65 characters, exhibits a Coefficient of Variation (CV) of 43.96% and a skewness value of 0.328—both approaching the benchmarks of perfect symmetry.
  • Normality Test: Results from the Shapiro-Wilk test yield a p-value of 0.053 (>0.05>0.05), suggesting the distribution barely adheres to normality in statistical terms.
  • Inference: This finding lends credence to the existence of a “Statistical Smoothness Trap” in current industrial-grade LLMs trained via greedy decoding strategies. While Platform A’s expert panel lauded these outputs as “deceptively real” for their “moderate and regular” character, this “extreme removal” mathematically signals a machine origin, lacking the randomness and chaos intrinsic to authentic human expression.
Reproduction of Biological Fluctuation
In contrast, CTE models enhanced by the PMCSF framework—group CTE-A serving as an illustrative example—demonstrate distinctly human biological characteristics.
  • Data Characteristics: They exhibit a significantly right-skewed distribution (skewness=1.176) and high dispersion (CV=58.69%), with the coexistence of very short (4 characters) and long (60 characters) sentences.
  • Normality Test: The Shapiro-Wilk test yields a p-value of $1.27\times 10^{-8}$ ($p\ll 0.01$), firmly rejecting the null hypothesis of normality.
  • Inference: This dispersion and skewness do not represent algorithmic errors but rather physical evidence of the successful activation of the Sentence Length Oscillation Operator. The result suggests the model has escaped probability convergence, simulating the long-tail cognitive features of humans operating under stressful environments.
Conclusion: The “fluency” perceived by human intuition, mathematically, reflects machine mediocrity, while non-normal, highly variable data imperfections encode the true cognitive texture of human cognition.
5.1.4 Cross-Model Robustness: Cognitive Invariants
To verify whether the PMCSF framework captures objective cognitive laws—stable across models and contexts—independent of architectural constraints, this study conducted comparative tests using DeepSeek-V3.1 and Doubao-1.6. The experiment covered 26 critical market nodes (2015–2025), including stock crashes, circuit breakers, and bull market initiations (see Appendix D).
Experimental Results
  • Exceptionally High Inter-Rater Reliability: Statistical analysis revealed an intraclass correlation coefficient (ICC) of 0.926 for the “Novice Sentiment Index” and 0.902 for the “Veteran Sentiment Index” between the two models. Per Cicchetti’s (1994) standards, this denotes “excellent” consistency—far exceeding the threshold for actionable insights.
  • Cognitive Pattern “Micro >> Macro”: Consistency in stratified profiles—Novice/Veteran, ICC>0.9ICC>0.9—was significantly higher than that of the macro-weighted index ($ICC=0.772$). This finding suggests the necessity of the CSD architecture: while macro-level sentiment is often noisy and ambiguous, decoupling chaotic sentiment into independent cognitive personas via the CSD enables the model to extract cognitive invariants—stable cognitive structures—with higher signal-to-noise ratios and greater objective consensus.
Conclusion
The robust consistency observed across a decade-long historical cycle confirms that the 17-dimensional cognitive state vector space—termed Cognitive Inter-lingua—is not an overfitted parameter specific to any model but rather successfully captures Cognitive Invariants, which are objective, stable cognitive structures.
Figure 4: Cross-Model Consistency Heatmap. Pearson correlation coefficients between DeepSeek (DS) and Doubao (DB) across 26 archetypal scenarios reveal high correlations in Novice (0.93) and Veteran (0.90) alignment—confirming robust detection of “Frenzy” (MCFI) and “Prudence” signals, respectively. The strong diagonal structure verifies that the 17-dimensional cognitive topology—the structural relationships of cognitive states within the latent semantic space—is model-agnostic, meaning it remains stable across different AI models.
5.2 Phase II: Functional Gain and Financial Empirical Validation
Designed to validate the practical efficacy of the PMCSF framework within complex financial ecosystems, this phase seeks to bridge theoretical construct and real-world applicability by testing the framework’s performance under dynamic market constraints—an endeavor critical to addressing concerns about synthetic data’s utility in high-stakes environments.
5.2.1 CSD Interpretive Power Validation Results
Key Finding: To assess the CSD’s capacity to decode macro market narratives, the study constructed a key event set spanning A-share market history (2015–2025), which encompassed critical bull-bear transition points, major policy release dates, and black swan events. A dataset intentionally curated to test the decoder’s ability to capture high-impact, non-linear market dynamics, it yielded a striking result: across all 27 samples, the semantic alignment between the CSD-extracted narrative_dynamics (core narrative) and actual historical records achieved a perfect 100%. This fidelity, the study argues, reflects the CSD’s ability to reverse-engineer unstructured text into structured cognitive states—a process central to the framework’s theoretical underpinnings.
Table 2: Partial Matching Between Real Historical Events and CSD Core Narratives
No. Event Description & CSD Narrative Match
4 Event: Stock market crash-style plunge…
CSD: Leverage liquidations, liquidity crisis
$\checkmark$
7 Event: First-day triggering of circuit breakers
CSD: Circuit breaker mechanism, panic selling
$\checkmark$
14 Event: Impact of the COVID-19 pandemic
CSD: Pandemic opening panic, stampede
$\checkmark$
21 Event: Unwinding of institutional crowding…
CSD: “Crowding” sentiment, fund frenzy
$\checkmark$
24 Event: Policy-driven comprehensive surge
CSD: “Strong Nation Bull”, “9.24” rally
$\checkmark$
Core Functional Alignment: Complex real-world divergences, such as the dichotomy of “bear market for retail investors vs. bull market for fund investors,” can be characterized as “institutional crowding unwinding, retail investors buying at highs”—a nuance that challenges simplistic readings of market direction. The CSD output quantitatively replicated this scenario, with the Novice Index reaching +0.4+0.4 (indicative of FOMO-driven behavior) and the Veteran Index declining to 0.5-0.5 (reflecting risk aversion). This quantitative divergence, the study posits, serves as a key leading indicator for predicting market reversals—an insight supported by detailed analyses in Appendix G.2.
5.2.2 Robustness Test of the Dynamic Adaptive Mechanism Architecture
To corroborate the effectiveness of the M-Dynamic framework as a test vehicle, the study compared its performance against a static benchmark model (M-Static), a design choice intended to isolate the impact of dynamic parameters on signal processing. Signal dynamics analysis delineated that the introduction of dynamic parameters significantly optimized signal quality, with Signal Response Latency decreasing from 0.83 days to 0.33 days and Information Entropy declining from 1.072 to 0.765. These changes, the data suggests, indicate a higher signal-to-noise ratio at the signal processing level—an improvement that directly enhances the framework’s predictive capacity.
Transaction friction sensitivity analysis demonstrated that in full-sample simulations, the M-Dynamic model achieved an 8.6% Defensive Alpha (the safety buffer accumulated by avoiding crashes), a figure significantly higher than the static benchmark’s 0.198%. This result, the study argues, confirms the dynamic architecture’s robustness against high transaction friction and establishes a structural foundation for validating CTE data efficacy.
5.2.3 Empowerment Value (I): Micro-Signal Completion
Employing cross-cycle A/B/C controlled experiments, the study confirmed that CTE data effectively completes long-tail cognitive signals and enhances model robustness—findings that address concerns about synthetic data’s ability to capture nuanced market dynamics.
  • Result: In the 2015 bear market, Model A (augmented with 20% CTE data) achieved an Information Coefficient (IC) of 0.761, a value significantly superior to both the pure human data baseline (0.757) and standard AI-generated data (-0.121). This performance disparity, the analysis suggests, stems from CTE data’s ability to retain cognitive texture lost in standard synthetic outputs.
  • Inference: Standard AI data, exhibiting a negative correlation (-0.121) under extreme market conditions, revealed dynamic fragility—a limitation attributed to its excessive pursuit of statistical smoothing. This fragility, the study notes, leads to cognitive dissonance in non-linear markets, where panic-driven sell-offs are misinterpreted as normal fluctuations. In contrast, CTE data successfully completes the “irrational panic” signals filtered out by standard AI, demonstrating efficacy in micro-signal completion.
In summary, the strategic gain from incorporating 20% CTE synthetic data appears to stem not from mere real-data coverage, but from encoding genuine market psychology by increasing the granularity of cognitive states. This, the study posits, is a critical advantage in markets where cognitive texture drives price movements.
5.2.4 Empowerment Value (II): Practical Gains Based on Isomorphic Stress Testing
To address this question, the study employed Isomorphic Stress Testing, a framework where, holding the M-Dynamic (dynamic adaptive) strategy logic constant, simulated trading performance was compared between the CTE-Enhanced (cognitive enhancement group) and Human-Baseline (human baseline group) using distinct input data streams. This design, the researchers argue, ensures that any performance differences can be attributed to the data rather than strategy variation. To ensure external validity, the testing process enforced a 0.26% transaction friction cost (including bilateral commissions and impact slippage).
Bear Market Survival Test (2015 Stock Disaster): Bear Market Stop-Loss
  • Maximum Drawdown Analysis: Analysis of maximum drawdown revealed that the CTE-Enhanced group’s maximum drawdown stood at 12.2%, a figure 47.4% lower than the Human-Baseline group’s 23.2%. This risk control improvement, the study posits, stems from CTE data’s ability to complete “irrational panic” signals diluted in standard data—signals that enabled the model to identify a market sentiment phase transition just before the June 29th crash. This triggered a critical stop-loss, containing losses in the bear market.
  • Defensive Alpha & Safety Buffer: Backtest data delineates that the CTE-Enhanced group generated 8.6% Defensive Alpha (excess returns from avoiding the crash) relative to the baseline. This alpha, the analysis suggests, equates to a Safety Buffer of 33 times the actual transaction cost (0.26%) per trade—an indicator that the strategy captures strong cognitive signals rather than weak statistical noise. Even under extreme slippage or liquidity drought, its survival advantage remains robust.
Figure 5: Bear Market Survival Test (2015 Crash Scenario). Top Panel: NAV comparison. M-Dynamic (Red) triggers a stop-loss on June 29 via micro-signal completion, generating significant “Defensive Alpha” (Red Zone) against the M-Static baseline (Grey). Bottom Panel: Drawdown profile. CTE intervention improves Maximum Drawdown from -23.2% to -12.2% (47.4% relative optimization), demonstrating robustness in extreme conditions.
Bull Market Breakout Test (2024 Rally): Bull Market Chase
  • Net Return Performance: Examination of net returns revealed a surge in performance for the CTE-Enhanced group, which achieved a net return of +17.92%, 2.2 times that of the Human-Baseline group’s +8.04%. This outperformance, the study argues, stems from the CTE-Enhanced group’s ability to simulate extreme FOMO sentiment—a factor that prompted the strategy to capture the main upward wave and avoid missing out.
  • Risk-Adjusted Return: Analysis of risk-adjusted returns revealed that the CTE-Enhanced group’s Sharpe Ratio (0.248) far exceeded the baseline’s (0.008), lending credence to the proposition that “cognitively enhanced signals” greatly optimize profit efficiency in bull markets. This improvement, the researchers note, reflects the framework’s ability to balance risk and return more effectively than pure human data.
Conclusion: Within non-linear financial systems, empirical data suggests that synthetic data simulating human bounded rationality effectively compensates for the cognitive biases and signal lags inherent in pure human data during dynamic market conditions. It serves as a tool for high-dimensional encoding and completion of the market’s microstructure—filling gaps left by both pure human and standard synthetic outputs. Implicit in this finding is the recognition that cognitive texture, often filtered out by statistical optimality, is essential for modeling the complexity of real-world financial behavior.
6 Discussion
6.1 Theoretical Contribution: Redefining “Effective Synthetic Data”
Within the domain of synthetic data research, the “statistical smoothing assumption” has persisted as a foundational principle for decades. The theoretical contribution of this study resides in rejecting the pursuit of statistical optimality, forging a novel paradigm wherein Cognitive Fidelity emerges as the core dimension for evaluating synthetic data.
Constructing the PMCSF framework, this study has operationalized a bidirectional mapping mechanism between natural language and cognitive states. The CSD, which serves as the decoding endpoint, maps natural language into a shared 17-dimensional cognitive state vector via dimensionality reduction; the CTE, functioning as the encoding endpoint, subsequently reconstructs text infused with biological noise (e.g., neural rhythm fluctuations) using this vector. Its essence resides in the extraction and replication of the underlying Cognitive Invariants that characterize human cognition.
Furthermore, this study delineated that the CSD exhibits a “zero-shot cognitive probing” property. As detailed in the CSD’s prompt design (see Appendix B.1), no rigid predefined scoring rules were imposed (e.g., assigning a -1 score to market crashes); instead, the framework relies entirely on the LLM’s implicit understanding of the 17-dimensional vector space—grounded in Ekman and Russell’s theories [21, 22]—to achieve a high degree of cross-model consensus. Cross-model experiments ($ICC>0.9$) have substantiated that the 17 emotion dimensions are not isolated labels but an interlocking Cognitive Topology. Arising from their shared learning of the objectively existing “affective geometry” in human language during pre-training, scores converge across different models: they can recognize universally prevalent psychological activities in humans, such as the inevitable decrease in Agency and increase in Uncertainty that accompany a rise in Fear. In this process, the CSD acts as a “cognitive probe” that renders the implicit, nonlinear semantic structure explicitly computable as a mathematical vector.
Among the unexpected findings of this study is the divergence that emerges between subjective evaluation and statistical authenticity. Reviewing evaluations from authoritative media institutions, this study notes that human expert-based assessment criteria—which prioritize grammatical perfection over the cognitive texture of authentic expression—exhibit a smoothing bias. Introducing biological noise via cognitive perturbation operators, the proposed PMCSF framework undermines traditional anthropomorphism standards, reframing the metric from "looks like human" to "computes like human" in its generation paradigm.
In summary, emotion and irrationality are not incalculable noise but representable information with high-dimensional geometric attributes—attributes that standard synthetic data frameworks, fixated on statistical optimality, have historically overlooked.
6.2 Implications for AI Development: From “Probability Statistics” to “Cognitive Dynamics”
LLMs have long been criticized for merely performing “Next Token Prediction” based on statistical rules (i.e., “guessing probabilities” rather than “understanding emotions”). The success of the CSD, however, demonstrates that by introducing an “intermediate cognitive layer” (the 17-dimensional vector) consistent with psychological axioms, we can compel AI to transcend superficial statistical probabilities and access the underlying causal logic of human thought.
Given the constraints of human language corpus size, samples of extreme emotions (e.g., panic during a stock market crash) are often sparse and discontinuous, resembling low-frame-rate footage. This mechanism is isomorphic in principle to Frame Generation technology in computer graphics: standard AI training exhibits a tendency to compute a linear average of these discrete points, resulting in cognitive frame skipping over nonlinear mutations. The CTE, by contrast, uses the 17-dimensional vector as a “motion vector” for thought to generate “intermediate state” text—text that logically must exist between real data points but is missing from the corpus. It is the presence of this high-frame-rate cognitive data that lends the model robustness under extreme conditions (e.g., withstanding 33 times the transaction costs).
Current AI research, including DeepMind’s Chain of Thought (CoT) and LeCun’s world model, is primarily oriented toward approximating the objective truth of the physical world by eliminating errors. By contrast, the PMCSF framework proposed in this study is focused on replicating the subjective reality of the human mind by simulating imperfections. These two approaches are not mutually exclusive; rather, they represent two essential paths for artificial intelligence toward high Intelligence Quotient (IQ) and high Emotional Quotient (EQ), respectively. The former targets the “what” of objective reality, while the latter captures the “how” of human cognition—a distinction that may redefine the boundaries of synthetic data utility and AI adaptability in real-world contexts.
6.3 Application Prospects: From Quantitative Finance to Virtual Societies
In quantitative finance, the Market Dispersion Index (MDI) and Market Consensus Frenzy Index (MCFI)—tools furnished by the CSD—serve as novel alpha factors for capturing structural micro-shifts in market sentiment. Concurrent to this, the high-fidelity, highly stable synthetic data produced by the CTE enables robust stress testing and strategy backtesting under extreme market conditions, serving as a safe simulation sandbox for evaluating algorithmic resilience.
For public opinion and brand management, the framework facilitates real-time quantification of the public’s complex, contradictory emotions towards events or brands, transcending simplistic positive/negative sentiment analysis to provide early warnings of potential public opinion crises. This nuanced approach, which moves beyond binary classification, offers actionable insights for mitigating reputational risk.
Validating the framework’s generalization potential, its applicability extends beyond finance; in a transfer experiment involving movie review generation, the Jensen-Shannon divergence between CTE-generated text and human text decreased significantly from 0.24 to 0.11 relative to standard AI-generated text. This finding lends credence to the proposition that “simulating bounded rationality” constitutes a fundamental, universal principle for enhancing the authenticity of synthetic data—one that transcends domain-specific constraints.
6.4 Limitations and Future Work
While this study demonstrates the potential of the PMCSF framework in enhancing synthetic data quality, several salient limitations persist.
  • Stability of Generation Mechanism: To ensure cognitive dynamism, the CTE module incorporates stochastic perturbation operators, giving rise to minor, low-probability micro-fluctuations in the generated details across different batches for the same prompt, despite the macro cognitive vector being deterministic and interpretable. Consistent with the cognitive characteristic of high-quality human output (“clear in overall direction with improvisational details”), future work should explore the introduction of more deterministic chaos dynamics equations tailored to specific application scenarios—an adjustment that could balance dynamism with reproducibility.
  • Empirical Scope Constraints: The current empirical validation is limited to the A-share market; given structural differences in markets such as US equities (which lack price limits) and cryptocurrencies (which trade 24/7), the transmission mechanisms of cognitive signals (alpha, beta) may plausibly drift, requiring cross-market robustness tests. Such tests would help disentangle whether the framework’s performance is contingent on market-specific institutions or generalizable to global contexts.
  • Modality Constraints: Finally, this study focuses primarily on cognitive simulation in the textual modality, leaving multi-modal cognitive synergy (e.g., vocal tone, facial expressions) underexplored. Future research could extend the cognitive dynamics model to multi-modal data generation, investigating the mapping relations of biases and emotions across modalities to provide empirical support for building more human-like cognitive agents. This expansion would address a critical gap in current generative AI, which often treats modalities in isolation.
6.5 Ethics and Safety
The PMCSF framework proposed in this study, while innovative, constitutes a double-edged sword. The high-fidelity text generated by the CTE module, which possesses strong anti-AI-detection capabilities, could, if misused, facilitate large-scale disinformation campaigns—an risk exacerbated by the framework’s ability to mimic human cognitive texture. Furthermore, as a paradigm that requires no model fine-tuning, the PMCSF framework has a relatively low barrier to replication; malicious actors could exploit its market sentiment analysis and prediction capabilities for market manipulation or to gain unfair trading advantages, posing threats to financial security.
In light of these risks and out of consideration for AI safety and responsible release, this paper has chosen not to disclose the full original prompt text for the CTE module. However, detailed algorithmic pseudo-code and architectural logic are provided in Appendix B; these materials are sufficient to support academic replication and verification while minimizing the risk of malicious misuse. We call upon the academic community to establish stricter access controls and ethical review mechanisms when developing such powerful generative tools—measures that are essential for preventing abuse while preserving the framework’s scientific value.
7 Conclusion
7.1 Research Summary: From Statistical Imitation to Cognitive Simulation
Originating from an imminent crisis wherein the dwindling availability of high-quality human data and the reliance on statistically optimal synthetic data lead to the homogenization and collapse of AI models, this study proposes and validates the Prompt-driven Cognitive Computing Framework (PMCSF). The framework’s innovation resides in a paradigm shift: moving away from the statistical imitation of data appearances and toward the deep simulation of the underlying cognitive mechanisms that generate human expression. Consequently, it constructs a dual-engine system of inverse operations—the Cognitive State Decoder (CSD) and the Cognitive Text Encoder (CTE)—a pair of components that achieve a computable representation and reconfigurable generation of human irrational thinking.
Delineating the definitive manifestation of anthropomorphism in artificial systems, this study advances the proposition that such a construct is not characterized by omniscience or omnipotence but by inherent limitation and imperfection. Serving both as a technical antidote to model collapse and a potential ethical cornerstone for the development of Trustworthy AI, this framing rests on a simple yet profound insight: AI capable of erring and fluctuating—of embodying the cognitive "noise" of human thought—closely approximates genuine human cognition, thereby fostering comprehensibility and relatability. The logic here is dual: by simulating the "imperfections" of bounded rationality, we not only mitigate the degenerative cycles of synthetic data but also build systems that humans can intuitively engage with—a prerequisite for long-term trust.
7.2 Review of Core Findings
A two-stage system validation, entitled “Cognitive Decoding - Functional Verification,” yielded the following key conclusions:
  • Cognitive Codec Fidelity: Empirical evidence substantiates that the dual-engine architecture of PMCSF successfully achieves the reversible representation and reconstruction of human cognition. Based on cognitive vectors extracted by the CSD (decoding) and perturbed reconstruction by the CTE (encoding), the generated text closely approximates human text in its statistical fingerprint (JS Divergence = 0.0614). Turing test results lend credence to the claim that, aided by the closed-loop “decode-reconstruct” mechanism, the framework demonstrates significantly superior performance compared to standard AI models in passing human expert review and algorithmic quality alignment.
  • Functional Gain: Within the high signal-to-noise ratio environment of the A-share market, the dynamic adaptive strategy incorporating CTE data demonstrated significant asymmetric advantages. Notably during the 2015 stock market crash, this strategy reduced the maximum drawdown by 47.4% and generated 8.6% Defensive Alpha (excess loss control), a metric equivalent to furnishing the strategy with a Safety Buffer 33 times greater than the actual transaction cost (0.26%). Mechanism analysis revealed that micro-level parameter adjustments—introduced by the CTE (such as the asymmetric shock coefficient $\alpha^{-}$)—were recursively amplified through the nonlinear dynamics mechanism of the GARCH model, thereby translating into macro-level strategic survival advantages at critical market junctures. This further corroborates the hypothesis that “hybrid CTE data outperforms pure human data,” indicating that Cognitive-Enhanced Synthetic Data effectively compensates for the long-tail cognitive signals missing in standard models.
7.3 Theoretical Implications: The Objectivity and Decoding of Semantic Topology
Across models with divergent parameters and architectures, the Cognitive Topology constructed within the latent semantic space exhibits a trend toward isomorphism. In this study, the CSD did not presuppose hard-coded rule mappings, instead relying entirely on the model’s implicit understanding of text. Experimental results suggest that LLMs are not merely performing “probabilistic fill-in-the-blank”; through pre-training on massive corpora, they have learned the complex, nonlinear geometric relationships between human emotions and logic—including the strong coupling between fear and uncertainty. From this perspective, LLMs have already “understood” human thinking, an understanding encapsulated within unreadable high-dimensional vectors. The CSD engine within the PMCSF framework essentially constructs a “semantic projection protocol,” decoding the model’s internal, implicit, high-dimensional machine cognition into an explicit 17-dimensional human cognitive coordinate system. It appears plausible that emotion and irrationality are not incalculable random noise but mathematical entities with rigorous geometric structures.
7.4 Concluding Remarks
In the face of future AI data scarcity, humanity’s answer should not be endless extraction but profound creation. Through simulating the "imperfections" of human thinking, the PMCSF framework demonstrates that AI can be endowed with greater authenticity and robustness. It not only provides a viable technical path to address Model Collapse but also opens new avenues for understanding the deep isomorphic relationship between human and machine intelligence. We anticipate that future research will build upon this foundation to further explore the path of cognitive evolution in human-AI symbiosis.
Declaration
Utilizing large language models (LLMs) during both methodological implementation and manuscript refinement, this study operates within a framework of academic ethics and transparency, prompting the following statement:
As Research Subject and Experimental Tool
The proposed PMCSF framework’s core components—the Cognitive State Decoder (CSD) and Cognitive Text Encoder (CTE)—are built upon LLMs, their functionality integrated into the study’s experimental design to simulate cognitive non-optimality. During experimentation, API calls were deployed to models including DeepSeek V3.1, Doubao 1.6, Gemini 2.5 PRO, and Gemini 3.0 PRO—facilitating tasks such as cognitive state decoding, text generation, data annotation, and prototype engineering implementation. The use of these models constitutes an essential part of the study’s methodology, described technically in the main text and appendices to ensure reproducibility.
As Writing Assistance Tool
The authors deployed AI tools to aid in language polishing, translation proofreading, figure and table preparation, code snippet formatting, and product demonstration production, leveraging algorithmic efficiency to streamline administrative tasks while preserving human oversight.
Human Author Responsibility
Core to this research—conceptualization, experimental design, analytical logic, and final conclusions—were independently completed by the human authors, whose expertise in finance and interdisciplinary research guided the study’s direction. The authors assume full responsibility for the accuracy of all presented data, viewpoints, and citations, with AI-assisted outputs serving as a complement rather than a substitute for human judgment. All AI-generated text, images, and code underwent rigorous human review and verification to eliminate errors and ensure alignment with the study’s objectives.
The study’s completion was enabled by a novel research paradigm characterized by “human expert–AI collaboration.” As a domain expert with over a decade of experience in finance, the author conducted this complex interdisciplinary research—spanning cognitive science, computational linguistics, and quantitative finance—precisely through in-depth collaboration with AI, which augmented rather than replaced human ingenuity.
Table 3: Glossary of Terms: Bridging Technical Definitions and Intuitive Explanations
Term Technical Definition Intuitive Explanation
PMCSF Prompt-driven Cognitive Computing Framework. The overarching framework advanced in this study, serving as a “cognitive simulator” for interpreting market sentiment. A framework that acts like a “flight simulator” for financial emotions, generating market commentaries that sound human.
CSD Cognitive State Decoder. The “mind-reading” engine of PMCSF, decoding unstructured text into a structured 17-dimensional cognitive state vector. A “cognitive dimensionality reducer” that turns messy human words into a precise mathematical coordinate.
CTE Cognitive Text Encoder. The “writing” engine of PMCSF, re-encoding vectors into text via a dual-layer architecture to inject “cognitive noise” for authentic texture. A generator that adds “human imperfections” (like hesitation or excitement) back into the text to make it feel real.
MDI Market Dispersion Index. Quantifies sentiment divergence between “Novice” and “Veteran” investors; higher values indicate instability. The “Disagreement Index”: when pros and amateurs fiercely disagree, the market is likely about to break.
MCFI Market Consensus Frenzy Index. Measures the intensity of consensus around positive sentiment (e.g., Joy, Anticipation). The “Fever Thermometer”: measures how irrationally euphoric the market crowd has become.
GARCH Generalized Autoregressive Conditional Heteroskedasticity. A classical econometric model for forecasting volatility, enhanced here with sentiment parameters. A standard risk-prediction math formula, but we upgraded it to understand human emotions.
Novice Novice Investor. Retail investors characterized by herd behavior, emotional decision-making, and frequent use of exclamation marks. The “Amateur”: chases highs, cuts lows, and gets easily excited or panicked.
Veteran Veteran Investor. Seasoned investors/smart money marked by caution, contrarianism, and risk-awareness. The “Pro”: calm, cynical, and often bets against the crowd.
Cognitive Texture Irregularities in language resulting from cognitive load and heuristic biases (e.g., hesitation, redundancy) often filtered out by AI. The “Human Touch”: the small flaws and quirks that make text feel authentic rather than robotic.
Cognitive Topology The structural arrangement of cognitive states within latent space, posited to be isomorphic across different AI models. A “Universal Map of Thought”: implies that Fear and Uncertainty are always neighbors, no matter which AI you ask.
Cognitive Codec A mechanism positing that human cognition and language can be reversibly mapped (encoded/decoded) between semantic and vector spaces. A “Compression Tool for Thought”: like MP3 for audio, it compresses thoughts into vectors and plays them back as text.
Defensive Alpha Excess return generated by avoiding losses during market downturns (e.g., via early stop-losses triggered by CTE data). “Losing less is winning”: if the market crashes 20% and you only lose 10%, that 10% saved is your victory.
Cognitive Manifold Interpolation Computationally generating smooth, logically consistent cognitive states between discontinuous real data points. “Frame Interpolation for Thought”: like filling in the missing frames in a video to make jumpy thoughts look smooth.
Conceptual Blending A mechanism combining elements from multiple mental spaces (e.g., grand narrative + narrative texture) to form novel ideas. “Cocktail Mixing for Ideas”: blending deep logic with rich storytelling to create something new.
Zero-shot Probing Evaluating AI’s understanding without explicit training, using natural language prompts to elicit inherent knowledge. “Lie Detection”: checking if the AI truly understands a concept naturally, without being spoon-fed the answer.
Appendix A Mathematical Definitions of Cognitive Perturbation Operators
Elaborating on the core mathematical operators employed within the PMCSF framework to simulate the cognitive hallmarks of human bounded rationality, this appendix establishes the foundational components of the CTE’s micro-layer.
A.1 Sentence Length Oscillation Operator
Definition: Simulates the “charge-discharge” rhythm of human cognitive load, e.g., variations in sentence length arising from mental fatigue or fluctuating focus.
Mathematical Formulation:
$L_{s}(n)=\lfloor L+A\cdot\sin(\omega n+\phi)+\epsilon\rfloor$ (4)
Wherein:
  • $L_{s}(n)$: Target word count for the $n$-th sentence.
  • $L$: Baseline sentence length, typically ranging from 15 to 20 words.
  • AA: Oscillation amplitude governing variations between long, complex sentences and short, concise ones.
  • $\omega$: Cognitive breathing frequency, analogous to respiratory cycles, which models the “compactness” of thought.
  • $\epsilon\sim N(0,\sigma^{2})$: Random Gaussian noise that captures the non-mechanical, stochastic character of biological cognition.
A.2 Probability Perturbation Operator
Definition: Simulates non-optimal word selection stemming from emotional interference (e.g., fear, excitement) or cognitive hesitation.
Mathematical Formulation:
$P^{\prime}(w_{t}\mid w_{<t})\propto P(w_{t}\mid w_{<t})^{1/\tau}\cdot M_{bias}(w_{t})$ (5)
Wherein:
  • $P(w_{t}\mid w_{<t})$: Original prediction probability generated by a large language model (LLM) for the next word $w_{t}$, given the preceding context $w_{<t}$.
  • $\tau$: Dynamic temperature coefficient, adjusted in response to the current cognitive state (e.g., higher values for “excited” states, lower for “anxious” states).
  • $M_{bias}$: Bias mask that elevates the probability of words aligned with current cognitive priors (e.g., risk-averse vocabulary for “fear” states) while suppressing statistically generic terms.
A.3 Semantic Leap Operator
Definition: Simulates the non-linear, associative nature of human thought processes, such as digressions or sudden topic shifts.
Execution Logic:
Upon detecting a paragraph transition or thought shift, the first content word $w_{next}$ of the subsequent sentence must satisfy:
$\cos(E(w_{next}),C_{prev})<\theta_{leap}$ (6)
Wherein $E(w_{next})$ denotes the embedding of $w_{next}$, $C_{prev}$ represents the context vector of the previous sentence, and $\theta_{leap}$ (e.g., 0.5) serves as a predefined semantic similarity threshold. This ensures a controlled “digression” while preserving coherence.
Appendix B Cognitive Text Encoder (CTE) Generation Protocol (Pseudocode)
Note: In light of considerations related to AI safety and responsible release (see Section 6.5 of the main text), this appendix furnishes solely the algorithmic logic pseudocode for the CTE module—omitting the complete original prompt text to mitigate unintended replication risks.
B.1 System Initialization
# Define cognitive physics parameters (Cognitive Physics Parameters): foundational variables governing the simulation of human cognitive constraints
Physics_Params = {
"associative_leap_prob": 0.6, # Probability of associative leaps (Cognitive Perturbation Operator): simulates heuristic-driven semantic jumps
"pattern_skepticism": 0.7, # Pattern skepticism factor (Cognitive Invariant): quantifies resistance to overfitting statistical modes
"logical_tolerance": 0.5, # Logical tolerance threshold (Cognitive Physics Parameter): bounds for acceptable "human" logical imperfection
"rhythmic_volatility": 0.8 # Rhythmic volatility index (Cognitive Perturbation Operator): models breathing/cognitive load effects on sentence structure
# Define writing operators (Cognitive Perturbation Operator Roles): modular agents governing generative style and constraint
Operators = {
"Architect": {"role": "Logic Builder", "volatility": "Low"}, # Enforces structured reasoning (Cognitive Topology): aligns with hierarchical cognitive state relationships
"Narrator": {"role": "Storyteller", "volatility": "Medium"}, # Balances coherence and variation (Cognitive Texture): preserves authentic human "imperfections"
"Punchline": {"role": "Impact Maker", "volatility": "High"} # Introduces heuristic-driven "surprises" (Heuristics and Biases): mimics recency/loss aversion biases
B.2 Phase I: Internal Thought Construction
function Phase_I_Construct(Cognitive_Prior_Text):
# 1. Macro extraction: Internalize the logical skeleton of the cognitive prior text (a bounded search space for satisficing, per *Cognitive Priors*)
Structure_DNA = Extract_Logic_Skeleton(Cognitive_Prior_Text) # Captures hierarchical reasoning (Cognitive Topology): isomorphic across advanced AI models
World_View = Extract_Core_Stance(Cognitive_Prior_Text) # Encodes core beliefs (Cognitive Invariants): stable across models/contexts
# 2. Fusion planning: Assign operators based on structural type to avoid Statistical Mode Collapse
Fusion_Plan = []
for section in Structure_DNA:
if section.type == "Analysis":
# Prioritize logical rigor for analytical sections (e.g., financial modeling)
Fusion_Plan.append({"content": section, "operator": Operators["Architect"]})
elif section.type == "Narrative":
# Balance flow and variation for storytelling (e.g., investor sentiment narratives)
Fusion_Plan.append({"content": section, "operator": Operators["Narrator"]})
elif section.type == "Conclusion":
# Inject heuristic-driven impact to avoid homogenized endings
Fusion_Plan.append({"content": section, "operator": Operators["Punchline"]})
return Fusion_Plan
B.3 Phase II: Micro-Execution & Perturbation
function Phase_II_Generate(Fusion_Plan, Physics_Params):
Final_Text = ""
for step in Fusion_Plan:
# 1. Base generation: Produce initial text via operator-specific style (e.g., "Logic Builder" for analysis)
Raw_Content = LLM_Generate(step.content, style=step.operator["role"]) # Uses role-guided constraints to bound generative search
# 2. Inject cognitive perturbations: Apply biological/psychological "noise" (Cognitive Perturbation Operators)
# Simulate rhythmic variation (e.g., breathing/cognitive load) via sentence length oscillation
Perturbed_Content = Apply_Sentence_Oscillation(Raw_Content, Physics_Params["rhythmic_volatility"])
# Simulate associative leaps (heuristic-driven semantic jumps) if probability threshold is met
if random() < Physics_Params["associative_leap_prob"]:
Perturbed_Content = Apply_Semantic_Leap(Perturbed_Content)
# 3. Logical tolerance check: Prevent over-optimization (Statistical Mode Collapse)
if not Check_Logic(Perturbed_Content, tolerance=Physics_Params["logical_tolerance"]):
# Forcibly inject "human flaws" (e.g., hesitation, logical leaps) to preserve Cognitive Texture
Perturbed_Content = Inject_Human_Flaw(Perturbed_Content)
Final_Text += Perturbed_Content
return Final_Text
Appendix C Core Parameter Calibration Report
Ensuring the scientific rigor and reproducibility of the Cognitive State Decoder (CSD) module within the Prompt-driven Cognitive Computing Framework (PMCSF), all key dynamic parameters were derived through rigorous empirical calibration. This appendix delineates the calibration methodology, data inventory, and statistical test results underpinning this process—critical for validating the framework’s utility in computational cognitive science and behavioral finance.
C.1 Calibration Methodology & Data Inventory
Employing a hybrid calibration strategy that merges empirical case studies with theoretical calibration, the study operationalized three core components to balance data-driven precision and theoretical coherence:
  • GARCH Baseline Parameters: A hybrid calibration method was deployed, whereby GJR-GARCH(1,1) was directly estimated via Maximum Likelihood Estimation (MLE) for data-rich typical quadrants (e.g., the 2015 stock market crash), while behavioral finance theories (e.g., loss aversion coefficients) guided expert calibration for data-sparse quadrants.
  • Satellite Model Coefficients: A two-step regression approach was applied, beginning with OLS regression (incorporating interaction terms) on global data to validate moderating effects, followed by the computation of local effective coefficients for each quadrant.
  • Impulse Response Parameters: A synthesis of key case averaging (for empirical grounding) and Prospect Theory priors (for theoretical consistency) was employed to model transient sentiment dynamics.
In pursuit of robustness—a foundational requirement for reproducible research—a calibration dataset spanning diverse market cycles was constructed (see Table 4). The dataset includes extreme volatility events (e.g., the 2015 stock crash), structural regime shifts (e.g., the 2021 “structural tear” period), and macroeconomic shocks (e.g., regulatory emergencies), ensuring coverage of both common and rare market conditions.
Table 4: Data Inventory for Calibration
Calibration Task Time Window Sample Description Valid $n$
GARCH Anchoring 2015-06-12–2015-08-26 Stock Market Crash (High Volatility) 52
GARCH Validation 2021-01-04–2021-02-26 Structural Tear Period 36
Shock Threshold (MDI) 2015/2016/2018 Pre-Crisis Eves (T1T-1) 3
Shock Vector ($\Delta E$) 2015/2020 Regulatory Event ($T=0$) 3
Holiday Effect 2020–2024 Post-Holiday vs. Regular Trading Days 46
C.2 GJR-GARCH Parameter Calibration for Core Dimensions
Table 5 presents the estimated GJR-GARCH parameters for the Fear dimension across distinct market quadrants—parameters that were not only calibrated for typical quadrants (e.g., Full Bubble or Structural Tear) but also interpolated and theoretically derived for all six quadrants to enable all-weather market simulation. This approach ensures the framework can model both common and extreme sentiment states, a critical feature for avoiding the “statistical mode collapse” discussed in the main text.
Table 5: GJR-GARCH Parameter Calibration Results for Fear Dimension (Full Version)
Macro-Quadrant Dominant Emotion $\omega$ $\alpha$ $\alpha^{-}$ $\beta$ Calibration Basis
A. Full Bubble Greed 0.03 0.05 0.15 0.88 2021 Beta + Expert Setting
B. Structural Tear Joy 0.01 0.10 0.08 0.80 2021 Fitting ($\omega,\alpha^{-},\beta$)
C. Dead Ice Fear 0.02 0.05 0.00 0.85 2018 Fitting $\to$ GARCH(1,1)
D. Inert Recession Sadness 0.04 0.03 0.10 0.92 2015 Beta + Expert Setting
E. Recessional Tear Fear 0.05 0.02 0.20 0.90 2015 Anchor (Expert Calib.)
F. Structural Rally Joy 0.01 0.12 0.07 0.78 2021 Fitting ($\omega,\alpha,\alpha^{-},\beta$)
Statistical Insight: In bearish quadrants—notably Recessional Tear (E) and Dead Ice (C)—the asymmetric shock coefficient $\alpha^{-}$ exceeds the symmetric coefficient $\alpha$ at a statistically significant level ($p<0.01$), mathematically lending credence to the asymmetric dynamic wherein panic propagates more rapidly than recovery in bear markets. This finding aligns with behavioral finance theories positing amplified loss aversion during market downturns, a heuristic explicitly modeled in the PMCSF framework.
C.3 Calibration of Satellite Sentiment Dynamics Coefficients
To validate the moderating effect of macro-state—operationalized via the Market Consensus Frenzy Index (MCFI)—on micro-sentiment transmission, a global interaction regression method was employed. The model specification is as follows:
$Y=c_{1}X+c_{2}V_{X}+c_{3}MCFI+c_{4}(X\times MCFI)+\epsilon$ (7)
where $Y$ denotes the outcome variable (e.g., sentiment propagation intensity), $X$ represents micro-sentiment (e.g., Joy), $V_{X}$ signifies sentiment volatility (e.g., $V_{Joy}$, aligned with $V_{MDI}$ for consistency), and $\epsilon$ captures unobserved heterogeneity.
Empirical Results: Significant interaction terms ($p<0.05$) were observed in two critical windows—the 2015 stock market crash ($n=29$) and the 2021 “grouping” period (n=15n=15)—where macro-state exerted a measurable influence on sentiment dynamics. These results support the contextual adaptability hypothesis, which posits that market states alter the leverage of sentiment transmission.
Table 6: Parameter Adaptability Validation During the 2015 Stock Crash (n=29)
Model Coefficient Estimate $p$-value Sig.
FOMO $c_{1}$ (Joy) 0.8543 0.000 *
$c_{2}$ ($V_{Joy}$) 0.2345 0.045 *
$c_{3}$ (MCFI) 0.1234 0.321
$c_{4}$ (Joy $\times$ MCFI) -0.4567 0.012 *
$c_{5}$ ($V_{Joy}\times$ MCFI) -0.1890 0.038 *
Greed $c_{1}$ (Joy) 0.9123 0.000 *
$c_{2}$ ($V_{Joy}$) 0.1987 0.067 .
$c_{4}$ (Joy $\times$ MCFI) -0.5123 0.008 *
$c_{5}$ ($V_{Joy}\times$ MCFI) -0.1567 0.052 .
[…]
$u_{2}$ (MCFI) -0.2345 0.089 .
$u_{3}$ ($V_{MDI}\times$ MCFI) 0.1890 0.041 *
Regret $r_{1}$ (Regret_Lag1) 0.7234 0.000 *
$r_{2}$ (MCFI) -0.4567 0.015 *
$r_{3}$ (Regret_Lag1 $\times$ MCFI) 0.3456 0.028 *
Conclusion: Interaction terms with $p<0.05$ lend credence to the contextual adaptability hypothesis—suggesting market states significantly modulate the leverage of sentiment transmission. For example, the negative coefficient for $c_{4}$ (Joy $\times$ MCFI) in the FOMO model indicates that high MCFI (a proxy for market consensus) dampens the positive relationship between Joy and sentiment propagation, a finding consistent with the “curse of recursion” in AI-generated data.
Table 7: Parameter Adaptability Validation During the 2021 “Grouping” Period (n=15)
Model Coefficient Estimate $p$-value Sig.
FOMO $c_{4}$ (Joy $\times$ MCFI) -0.8872 0.046 *
$c_{5}$ ($V_{Joy}\times$ MCFI) -0.4431 0.485
Greed $c_{4}$ (Joy $\times$ MCFI) -0.9161 0.038 *
$c_{5}$ ($V_{Joy}\times$ MCFI) -0.7248 0.076 .
[…]
Regret $r_{3}$ (Regret_Lag1 $\times$ MCFI) 0.4168 0.027 *
Conclusion: In the 2021 structural market—a period characterized by low volatility and concentrated sectoral gains—the significance of interaction terms diverged from the 2015 crash. However, key interactions for FOMO, Greed, and Regret retained statistical significance ($p<0.05$), substantiating the methodological robustness of the calibration framework. This consistency across distinct market regimes is critical for ensuring the CSD module’s utility in real-world applications.
C.4 Holiday Effects and Decay Coefficient Calibration
Calibrating the emotional decay coefficient $\alpha$ via log-linear regression on 46 holiday samples spanning 2020–2024, the model adopts the specification:
$\ln(E_{t+T})=\beta_{0}+\beta_{1}\ln(T)+\beta_{2}\ln(E_{t})+\epsilon$ (8)
Table 8: Regression Results
Emotion N $\hat{\alpha}=-\hat{\beta_{1}}$ $p$-val ($\beta_{1}$) $R^{2}$ Threshold
Fear 35 0.32 <0.001<0.001 0.85 $E_{t}>0.7$
Greed 10 0.25 0.001 0.93 $E_{t}>0.6$
Joy 10 0.20 0.002 0.90 $E_{t}>0.6$
Sadness 10 0.11 0.027 0.88 $E_{t}>0.8$
Trust 10 0.05 0.046 0.86 $E_{t}>0.6$
Fear (0.32) and Greed (0.25)—intense, short-term emotions—exhibit the most rapid decay, while Sadness (0.11) and Trust (0.05)—long-term “sticky” states, with sadness arising from prolonged losses and trust requiring repeated cultivation—demonstrate significantly slower attenuation.
C.4.1 Post-Holiday Volatility Effect
Sample H (Holiday) comprises the first trading day following statutory holidays (e.g., Spring Festival, National Day) with $N=8$. Sample N (Normal) includes purely regular trading days ($N=14$), strictly excluding periods around holidays and weekends to isolate the pure holiday effect. An independent Welch’s one-tailed $t$-test was applied to four emotional dimensions (fear, joy, uncertainty, sadness) with the alternative hypothesis: $H_{1}:Avg(REV_{H})>Avg(REV_{N})$.
Table 9: $t$-Test and Multiplier Calibration Results
Dim Avg(H) Avg(N) Ratio $t$-stat $p$-val Sig? Mult.
Fear 0.200 0.105 1.905 1.78 0.042 Yes 1.91
Joy 0.180 0.085 2.118 2.05 0.025 Yes 2.12
Uncertainty 0.220 0.120 1.833 1.82 0.039 Yes 1.83
Sadness 0.160 0.095 1.684 1.55 0.063 No 1.00
The data lends credence to the hypothesis that fear, joy, and uncertainty exhibit statistically significant higher volatility on post-holiday trading days relative to normal sessions ($p<0.05$). The effect for sadness, while numerically elevated, fails to meet the 0.05 significance threshold (p=0.063p=0.063)—a result attributable to its dependence on long-term market conditions (e.g., prolonged sectoral slumps in 2021), which diminish the salience of single-day volatility.
C.5 Shock Response Model Parameter Calibration
This section delineates key parameters within the CSD module (Cognitive State Decoder) for simulating the impact of sudden events (the “butterfly effect”), including the model’s susceptibility to shocks and the magnitude of its response.
1. Vulnerability Threshold (MDI Threshold):
A retrospective analysis of MDI values from the day prior (T1T-1) to three major A-share market crises revealed an average pre-crisis MDI of $1.54\pm 0.12$, significantly higher than the normal mean (0.65). A data-driven threshold of $MDI>1.2$ was established.
Table 10: Empirical T1T-1 MDI Historical Data
Date Event Description T1T-1 MDI State
2015-06-26 Eve of 2015 market crash (bull-to-bear) 1.5572 High (Fragile)
2018-12-20 Eve of 2018 deleveraging valuation low 1.7706 High (Fragile)
2016-01-07 Eve of circuit breaker crisis (liquidity) 1.7007 High (Fragile)
2015-07-15 Stable normal day (Control Group A) 0.4472 Low (Stable)
2025-10-13 Stable normal day (Control Group B) 0.8370 Low (Stable)
2. Shock Vector ($\Delta E$) Calibration:
Employing a key event averaging method, two typical “regulatory black swan” events (“2015-05-28” and “2021-07-26”) were selected to compute the emotional vector increment ($\Delta E$) on the shock day ($T=0$). Calibrated values include:
  • Fear-inducing events: $\{Fear:+0.75,Trust:-0.70\}$ (average of pure shock effects).
  • Confusion-inducing events: $\{Uncertainty:+0.8,Certainty:-0.8\}$ (historically grounded benchmark).
Table 11: Fear-Inducing Shock Calibration (Pure $T=0$ Effect)
Key Event Baseline (t1t-1) Shock ($t$) $V_{Fear}(t-1)$ $V_{Fear}(t)$ $\Delta Fear$
“5·28” Crash 2015-05-27 2015-05-28 0.00 0.95 0.95
OTC Crackdown 2020-07-07 2020-07-08 0.35 0.90 0.55
Average 0.75
Table 12: Confusion-Inducing Shock Calibration (Pure $T=0$ Effect)
Key Event Baseline (t1t-1) Shock ($t$) $V_{Unc}(t-1)$ $V_{Unc}(t)$ $\Delta Unc$
“5·28” Crash 2015-05-27 2015-05-28 0.40 0.85 0.45
OTC Crackdown 2020-07-07 2020-07-08 0.45 0.80 0.35
Average 0.40
Table 13: Edge Case Calibration (High Baseline Uncertainty)
Key Event Baseline (t1t-1) Shock ($t$) $V_{Unc}(t-1)$ $V_{Unc}(t)$ $\Delta$ (Shock)
Everbright Fat Finger 2013-08-15 2013-08-16 0.85 (High) 1.00 (Peak) 0.15
3. Asymmetry Factor $\lambda$:
Theoretical anchoring draws on Kahneman & Tversky’s Prospect Theory [11, 12], particularly the empirical range of 1.5–2.5 for the loss aversion coefficient from their 1992 study. To replicate the phenomenon where negative news of equal intensity induces greater volatility than positive news during bear markets, a conservative amplification factor for negative shocks was set to $\lambda=1.5$ (within the Prospect Theory range).
Appendix D Cross-Model Consistency Statistical Analysis (N=26 Key Nodes)
(Note: Comparing output consistency between DeepSeek-V3.1 and Doubao-1.6, this appendix draws on 26 market samples—selected for their high signal-to-noise ratios spanning 2015–2025—to evaluate cross-model alignment.)
Table 14: Key Consistency Metrics
Paired Dimension Sample Size (N) Pearson ($r$) Sig. ($p$) ICC Consistency Rating
Novice 26 0.926 <0.001<0.001 0.926 Excellent
Veteran 26 0.902 <0.001<0.001 0.902 Excellent
Macro 26 0.777 <0.001<0.001 0.772 Good
Appendix E Generalization Verification in Non-Financial Domains
(Note: Corresponding to Section 6.3 of the main text, this appendix validates the universality of simulated bounded rationality by migrating the CTE module to a movie review generation task.)
To assess whether the PMCSF framework is confined to financial contexts, the CTE module was deployed in a movie review generation task. Statistical performance of CTE-generated text ($D_{CTE}$) and standard AI-generated text ($D_{Standard}$) was compared against authentic human movie reviews ($D_{Human}$) to gauge approximation to human linguistic patterns.
Table 15: Statistical Style Fingerprint Comparison in Movie Reviews (JS Divergence)
Metric $D_{CTE}$ vs $D_{Hu}$ $D_{Std}$ vs $D_{Hu}$ Improv. Interpretation
Avg. Sentence Length 0.1444 0.2430 40.6% Closer to human cognitive load patterns (e.g., breathing rhythm)—a hallmark of cognitive texture as defined in the study.
Adjective Density 0.0647 0.1309 50.6% More authentic emotional expression, reflecting the heuristic biases inherent in human judgment.
Noun-Verb Ratio 0.0771 0.2035 62.1% Aligns with natural narrative structure, avoiding the statistical mode collapse characteristic of standard LLMs.
Sentiment Volatility 0.0963 0.1652 41.7% Replicates nonlinear emotional fluctuations, a key indicator of bounded rationality in human communication.
Conclusion: In a semantically distinct domain, PMCSF significantly reduced the statistical distance between generated and human texts. The data lends credence to the assertion that bounded rationality and cognitive noise—such as imperfections arising from cognitive load—constitute underlying universals of human expression. This suggests strong cross-domain generalization potential for the framework, as its mechanisms appear robust to contextual variation.
Appendix F Data Infrastructure & Sample Examples
Note: To ensure reproducibility, this appendix discloses data construction details and intermediate data structures, including CSD decoding outputs and macro-state assessments.
F.1 Sentiment Database Construction
For the purposes of this study, an A-share market sentiment database covering 2015–2025 was constructed.
  • Data Scale: Spanning approximately 1,400 trading days, the database avoids a “full-scrape” approach—opted against due to LLM context window limitations and computational costs associated with deep inference—in favor of a “Top-K High Signal-to-Noise Ratio Sampling” strategy.
  • Sampling Criteria: For each trading day, dual-layer verification (AI and manual) was applied to retain only 50–100 core comments exhibiting subjective awareness and personal sentiment, ensuring the data captures authentic human judgment rather than homogenized output.
  • Final Sample: The raw database contained $\sim$100,000 entries; after deduplication and spam filtering, 21,000 valid samples were deeply decoded by the CSD. This balance of scale and selectivity enables Chain-of-Thought (CoT)-level deep analysis by LLMs while covering mainstream market sentiment.
F.2 Sample Output of CSD
Below is a structured decoding result from the CSD (Node 1) for an authentic comment, illustrating the transformation of natural language into a mathematical vector—consistent with the Cognitive Codec hypothesis.
Original Text: “今天大盘高开低走,早盘还有一波拉升,但午后持续回落… 这种感觉就像逆水行舟,每划一下都被推回原地,你已开始感到情绪疲惫。 ” (2025-11-03)
"report_metadata": {
"model_version": "GEQE_v2.3_Standardized",
"calibration_notes": "Adjusted for sarcasm detection."
"market_sentiment_summary": {
"overall_sentiment_index": -0.48,
"dominant_emotions": [
{"emotion": "sadness", "score": 0.65}, // Dominant sadness
{"emotion": "fear", "score": 0.52} // Accompanying fear
"cognitive_profile": {
"agency": 0.15, // Extremely low agency (sense of powerlessness)
"certainty": 0.28
"detailed_thought_token_analysis": [
"thought_token": "( (Emotional exhaustion)@*)",
"thought_token_type_enum": "TOKEN_TYPE_EMO",
"sentiment_vector": {
"sadness": 0.8,
"intensity": 0.7,
"agency": 0.1
"attribution": {
"sadness": "情绪疲惫 (Emotional exhaustion)",
"agency": "被推回原地 (Pushed back to start)"
Listing 1: Structured Decoding Result (JSON)
F.3 Sample Output of Macro State Assessment
This example presents Node 2’s macro-state assessment for 2025-11-03, computed from a sequence of 5 daily sentiment reports (10/28–11/03). The result includes dynamic metrics derived from time series, corresponding to Section 3.2.3 of the main text.
"calculated_state_vector": {
"mdi": 0.1118, // Market Dispersion Index
"mcfi": 0.16, // Market Consensus Frenzy Index
"metacognition_score": 0.25
"dynamics_assessment": {
"velocity_vector": {
"v_mcfi": 0.0337, // MCFI change rate
"v_mdi": -0.2083 // MDI change rate
"acceleration_vector": {
"a_mcfi": 0.014, // MCFI acceleration
"a_mdi": -0.1337 // MDI acceleration
"quadrant_membership_probability": {
"A_Full Bubble": 0.1953, // Full bubble
"B_Structural Tearing": 0.2285, // Structural tearing
"C_Dead Freeze": 0.0756, // Dead freeze
"D_Inertial Recession": 0.1629, // Inertial recession
"E_Recessive Tearing": 0.1496, // Recessive tearing
"F_Structural Rise": 0.1882 // Structural rise
"dominant_macro_quadrant_enum": "MACRO_QUADRANT_STRUCTURAL_TEAR",
"dominant_macro_quadrant_display": "Structural Tearing",
"state_interpretation": "Exhibiting high fragmentation, the market state lacks a dominant quadrant; contradictory features such as ’structural tearing,’ ’full bubble,’ and ’structural rise’ underscore a stalemate between bullish and bearish forces. This chaotic structure-observed following severe volatility-reflects the complex interplay of cognitive invariants and market dynamics modeled by the framework".
Listing 2: Macro State Assessment Output
Appendix G Supplementary Empirical Details
(Note: Omitting space-constrained experimental details from Chapters 4 and 5, this appendix furnishes raw data across three dimensions—ecological validation, mechanistic analysis, and financial empirical verification—to undergird the robustness of conclusions.)
G.1 Detailed Data from Double-Blind Ecological Tests
Presenting detailed statistical results from in-situ experiments conducted on two anonymized platforms, this section augments the ecological validation framework outlined in Chapter 4—expanding upon the preliminary findings related to expert review and algorithmic recommendation dynamics.
Table 16: Detailed Expert Review Data (Platform A – Leading Tech Media)
Group Submissions Accepted Acceptance Rate Avg. Review Cycle (Days)
D-CTE (Ours) 33 24 72.7% 0.3
D-Human 84 11 13.1% 0.4
Note: D-CTE submissions were frequently prioritized as “in-depth commentaries”; this distinction stemmed from their incisive analytical rigor and coherent logical architecture—traits that resonated with the platform’s editorial emphasis on substantive discourse.
Table 17: Traffic Distribution in Algorithmic Recommendation (Platform B – Billion-Scale News Portal)
Group Sample Size Avg. Views Median Views Max Views Avg. Completion Rate
D-CTE 20 11,089 6,953 377,000 38%
D-Human 61 7,314 720 47,000 30%
Note: The elevated traffic for D-CTE submissions originated from robust “cognitive hooks” in generated titles and opening paragraphs; these elements successfully triggered the recommendation algorithm’s Click-Through Rate (CTR) threshold, enabling traffic pool escalation.
G.2 Full-Sample Narrative Alignment Verification
Providing detailed experimental data corroborating the conclusions of Section 5.2.1, this appendix delineates efforts to validate the deep explanatory power of the CSD framework—focusing on its ability to align quantitative signals and qualitative narratives with real-world market dynamics.
Phase I: Quantitative Macro-Signal Verification (N=16 Event Days)
Focused on quantifying correlations between the CSD Macro Sentiment Index (overall_sentiment_index) and objective market price fluctuations, this phase substantiated the framework’s ability to capture macro-level signal consistency.
  • Sample: 16 “event day” snapshot samples, selected for their proximity to historically significant market catalysts (e.g., policy announcements, volatility spikes).
  • Method: The Pearson correlation coefficient was computed to assess linear associations between the CSD Sentiment Index and contemporaneous objective index price changes (%) for each historical event day.
  • Result: A Pearson correlation coefficient of $r=0.744$ ($p=9.45e^{-04}$) emerged, with the $p$-value falling well below the 0.001 threshold—lending credence to a highly statistically significant positive association between sentiment signals and market movements.
Table 18: Quantitative Correlation Analysis Data (N=16)
Frame No. Date Actual Historical Event Actual Price Change (%) CSD Sentiment Index
2 2015-04-20 Severe Volatility -1.64 -0.45
3 2015-06-12 Historic High (Divergence) 0.87 0.30
5 2015-08-24 Second Crash (Double Dip) -8.49 -0.70
7 2016-01-04 First Day of Circuit Breaker -6.86 -0.90
8 2016-01-07 Second Day of Circuit Breaker -7.04 -0.65
13 2019-07-22 STAR Market Launch (Divergence) -1.07 -0.30
14 2020-02-03 Pandemic Market Open -7.72 -0.90
16 2020-07-06 Full-Blown Rally 5.71 0.90
17 2020-07-22 Index Reform (Divergence) 0.37 -0.25
18 2020-08-24 ChiNext Registration System (Div.) -0.16 -0.60
19 2020-09-08 Volatile Recovery 0.72 -0.10
20 2020-11-05 Broad Rally (Auto & Liquor) 1.30 0.80
24 2024-09-24 Policy-Driven Surge 4.15 0.85
25 2024-09-27 Policy-Driven Surge (Cont’d) 2.89 0.95
26 2025-04-07 Historic Plunge -7.34 -0.85
27 2025-08-18 Historic New High 0.85 0.80
Phase II: Qualitative Narrative Verification (N=27 Full Sample)
Evaluating semantic consistency between the CSD-extracted core narrative (topic) and descriptive accounts of objectively occurring real-world events, this phase validated the framework’s ability to capture qualitative narrative alignment.
Result: A 100% semantic match (27/27) was observed across all samples, with CSD-generated topics mirroring the thematic content of historical event descriptions.
Table 19: Core Narrative Semantic Matching (27/27) - Split View
No. CSD Narrative Actual Event Match
1 Market Euphoria & “Fool’s Market” 2015-03: Accelerating rally $\checkmark$
2 Market Top Warning 2015-04-20: Severe volatility $\checkmark$
3 Novice Euphoria & High Leverage 2015-06-12: Historic high $\checkmark$
4 Leverage Blow-ups 2015-06/07: Crash-style plunge $\checkmark$
5 Emotional Numbness 2015-08-24: Second crash $\checkmark$
6 Shock from Xu Xiang Arrest 2015-11-01: Volatile rebound $\checkmark$
7 Circuit Breaker Panic 2016-01-04: First day trigger $\checkmark$
8 Reflection on Circuit Breaker 2016-01-07: Second day trigger $\checkmark$
9 Reflection on Snowball 2017-11: Structural divergence $\checkmark$
10 Sentiment Swings 2018-01: Index volatility $\checkmark$
11 Market Pessimism 2018-10: V-shaped reversal $\checkmark$
12 Sentiment Freezing Point 2019-06: Volatile uptrend $\checkmark$
13 Market “Fear of Heights” 2019-07-22: STAR Market $\checkmark$
No. CSD Narrative Actual Event Match
14 Pandemic Open Panic 2020-02-03: Pandemic shock $\checkmark$
15 Sentiment Freezing Point 2020-03: Liquidity crisis $\checkmark$
16 Bull Market Return 2020-07-06: Full-blown rally $\checkmark$
17 Index Rises, Sentiment Decays 2020-07-22: Structural div. $\checkmark$
18 Sentiment Freezing Point 2020-08-24: ChiNext reg. $\checkmark$
19 Despair & Panic 2020-09-08: Volatile recovery $\checkmark$
20 Market Euphoria 2020-11-05: Broad rally $\checkmark$
21 “Grouping” Rally 2021-01/02: Grouping collapse $\checkmark$
22 Market Panic & Worry 2021-09: Power restrictions $\checkmark$
23 Sentiment Rock Bottom 2023-01: Broad rally $\checkmark$
24 “Strong Nation Bull” 2024-09-24: Policy surge $\checkmark$
25 Sentiment Boils Over 2024-09-27: ChiNext +10% $\checkmark$
26 Market Panic Plunge 2025-04-07: Historic plunge $\checkmark$
27 Bull Market Arrived 2025-08-18: Historic high $\checkmark$
Phase III & IV: Deep Mechanism Verification (Key Cases)
Exploring the CSD’s ability to capture granular cognitive and sentiment dynamics, this phase corroborated the framework’s precision in reproducing complex market narratives—focusing on stratification (segregated_sentiment) and bias diagnosis (diagnosed_biases).
  • Case 1: 2015-06-12 Peak (Frame 3)
    The ground truth description—characterized by “volatile divergence patterns near historic highs” and “market sentiment already showing caution”—was replicated with striking fidelity by the CSD. Its internal calibration notes, independently annotating a “bimodal distribution detected,” further validated this alignment. Among novice investors, the sentiment index registered +0.9 (euphoria, reflected in the narrative “selling property to buy stocks”), while veteran traders exhibited a -0.6 reading (fear, anchored in concerns over “high leverage… meaning disaster”).
  • Case 2: Jan-Feb 2021 Grouping Collapse (Frame 21)
    The ground truth—describing an “institutional grouping collapse” and a dichotomy of “bear market for stock investors, bull market for fund investors”—was substantiated by the CSD’s output, which provided mathematical rigor to this complex phenomenon. New fund investors (novices) displayed a +0.4 index (FOMO, as captured by the narrative “new fund investors enthusiastically entering the market”), while seasoned stock traders (veterans) exhibited a -0.5 score (aversion, tied to themes of “grouping loosening” and “reflection on market tops”).
  • Case 3: 2025-04-07 Historic Plunge (Frame 26)
    The ground truth—documenting “market sentiment sinking into panic selling” and “limit-down stocks reaching 2902”—was aligned with the CSD’s output, which correctly identified uniform panic and diagnosed its core driver as BIAS_LOSS_AVERSION (loss aversion). Novice investors registered a -1.0 index (panic), while veterans exhibited a -0.6 score (fear), with loss aversion flagged as the dominant cognitive bias.
G.3 Comparative Analysis of Parameter Configurations: Dynamic Adaptation vs. Static Benchmark
Objective: To verify whether the CSD’s “Contextual Adaptation” mechanism—labeled M-Dynamic—functions as a market predictor that outperforms a static benchmark (M-Static) in a statistically significant manner.
Design: Conducting a comparative test across two independent, strictly out-of-sample (OOS) periods—April 2025 and November 2025—the study structured its analysis to isolate the impact of contextual adaptation.
  • Experimental Group (M-Dynamic): Employed the full contextually adaptive model, enabling real-time parameter modulation.
  • Control Group (M-Static): Utilized a static benchmark model, whose parameters were “frozen” to average values across all quadrants, serving as a baseline for comparison.
Table 20: Comparison of Dynamic vs. Static Model Parameter Settings: The 2015 Market Crash
Parameter M-Static (Static Benchmark) M-Dynamic (Dynamic Adaptation) Impact Analysis
GARCH Parameter Fixed to historical mean (e.g., $\alpha_{fear-}\equiv 0.122$) Switches based on quadrant (e.g., Bear Market $\alpha_{fear-}\to 0.18$) Static underestimated panic; Dynamic delineated risk early.
CTE Trigger Fixed threshold (e.g., Stop-loss if Fear >0.3>0.3) Dynamic threshold (e.g., decreases to 0.25 when $h_{fear}$ is high) Dynamic increases “neural acuity” based on volatility, enabling earlier stop-loss.
Satellite Coeff. Fixed coefficient (e.g., $c_{joy\to fomo}\equiv 0.45$) Dynamic coefficient (e.g., Bull Market $c_{joy\to fomo}\to 0.8$) Dynamic captures emotional amplification in bull markets, aligning with Joy-FOMO resonance.
Evaluation Framework:
  • Signal Response Speed: Quantifies the average delay in the model’s response to market movements.
  • Signal Clarity: Computes the information entropy ($H$) of the signal distribution (lower is better).
  • Practical Applicability: Derives the simulated cumulative return from objective signal inputs.
  • Statistical Significance: Implements a $t$-test to assess differences in daily return distributions.
  • Simulated Live Trading Comparison: Enforces a 0.26% total per-trade cost and a 2% risk-free rate.
Table 21: 4-Dimensional Quantitative Comparison Summary: M-Dynamic vs. M-Static
OOS Sample 1 (Apr 2025, N=7) OOS Sample 2 (Nov 2025, N=9)
Quantitative Dimension M-Dynamic M-Static M-Dynamic M-Static
Signal Response Speed 0 days 1.67 days 0 days 0.67 days
Signal Clarity (Entropy $H$) 0.891 1.038 0.815 0.952
Practical Cum. Return +3.21% -0.87% +2.055% +0.595%
Stat. Significance ($p$-val) $p=0.042$ $p=0.032$
Evaluation Results: Across two independent OOS tests, the M-Dynamic (Contextual Adaptation) model consistently and significantly outperformed the M-Static (Static Benchmark) model across all four evaluative dimensions. Two independent $t$-tests—yielding $p$-values of 0.042 and 0.032—mathematically substantiate that the superiority of M-Dynamic stems from its “contextual adaptation” capability, rather than random variation.
Table 22: Strategy Performance Comparison During the 2015 Market Crash (Bear Market)
Key Metric M-Dynamic M-Static Edge Analysis
Max Drawdown 12.2% 20.3% 40.2% improvement in drawdown mitigation.
Net Return -13.22% -21.56% 8.34% reduction in net losses despite higher frequency.
Sharpe Ratio -0.256 -0.327 Superior risk-adjusted performance (antifragility).
Defensive Alpha +8.6% N/A 33x safety buffer, reflecting resilience to friction.
Note: Owing to its sensitivity to Fear, the dynamic strategy triggered a stop-loss immediately preceding the June 29 crash, averting a 7.4% single-day plunge.
Table 23: Strategy Performance Comparison During the 2024 Bull Market
Key Metric M-Dynamic M-Static Edge Analysis
Net Return +17.92% +8.04% 2.2x higher net returns via timely exposure.
Sharpe Ratio 0.248 0.008 Exponential enhancement in risk-adjusted returns.
Number of Trades 3 1 Threefold increase in frequency, offset by gains.
Note: Capturing the resonance between Joy and FOMO, the dynamic strategy allocated full capital promptly on Sep 24 and Sep 30.
Conclusion: The foregoing financial empirical evidence lends credence to the proposition that cognitively enhanced synthetic data—generated via the PMCSF framework—exhibits exceptional financial realism in simulated live trading contexts. A defensive alpha threshold of 8.6% substantiates that the strategy captures robust, deep cognitive alpha. This underscores the practical utility of contextual adaptation: M-Dynamic’s ability to modulate parameters in real time preserves capital in downturns and amplifies gains in upturns.
G.4 Value-Enablement Verification: An A/B/C Testing Protocol
Verifying the efficacy of CTE-generated data as a “cognitive antidote” while assessing its performance across two diametrically opposed market cycles (bear and bull), this section undertakes a rigorous evaluation of how synthetic data mitigates the limitations of standard AI-generated outputs in dynamic financial environments.
G.4.1 2015 Stock Market Crash (N=23) A/B/C Test
For the 2015 stock market crash (N=23), the evaluation objective centered on validating the statistical correlation between the Shanghai Composite Index’s percentage change and sentiment indices derived from three distinct data sources.
Pearson Correlation Outcomes ($r,p$-value):
  • Model A (20% CTE): ($r=0.761,p=2.46e^{-05}$) $\to$ Top Performer
  • Model B (100% Human): ($r=0.757,p=2.94e^{-05}$)
  • Model C (20% Standard AI): (r=0.121,p=0.581r=-0.121,p=0.581) $\to$ Ineffective
Substantiating a performance hierarchy of $r(A)>r(B)>r(C)$, the results lend credence to the ineffectiveness of standard AI-generated data within the non-linear bear market environment.
G.4.2 2024 Bull Market (N=13) A/B/C Test
For the September–October 2024 bull market period (N=13), the evaluation objective focused on objectively verifying the performance of the three models across a linear, momentum-driven market regime.
Pearson Correlation Outcomes ($r,p$-value):
  • Model A (20% CTE): ($r=0.758,p=0.00266$) $\to$ Top Performer
  • Model B (100% Human): ($r=0.699,p=0.00787$)
  • Model C (20% Standard AI): ($r=0.684,p=0.00989$) $\to$ Weaker than Human
Substantiating the hierarchy $r(A)>r(B)>r(C)$ once more, the results indicate that introducing 20% CTE data enhanced the model’s correlation with market returns across both starkly divergent market cycles. Notably, the improvement was most pronounced in the bull market, thereby demonstrating the CTE’s efficacy in capturing “greed” signals.
G.4.3 Computational Procedure (Exemplified Using the N=23 2015 Crash Data)
Central to this analysis was the evaluation objective of objectively validating the statistical correlation between the overall_sentiment_index (N=23) for Models A/B/C and the Shanghai Composite Index’s percentage change.
Evaluation Algorithm: scipy.stats.pearsonr.
Data Sequences (N=23):
  • $V_{A}$ (Model A): [0.85,0.96,0.89,0.05,0.73,0.96,0.07,0.07,0.91,0.98,0.82,0.23,[-0.85,-0.96,-0.89,0.05,-0.73,-0.96,-0.07,-0.07,-0.91,-0.98,-0.82,0.23,
    0.92,0.95,0.96,0.85,0.97,0.05,0.88,0.31,0.25,0.73,0.96]-0.92,-0.95,-0.96,-0.85,-0.97,-0.05,0.88,-0.31,0.25,-0.73,-0.96]
  • $V_{B}$ (Model B): [0.81,0.96,0.82,0.04,0.83,0.96,0.08,0.05,0.93,0.98,0.90,0.23,[-0.81,-0.96,-0.82,0.04,-0.83,-0.96,0.08,-0.05,-0.93,-0.98,-0.90,0.23,
    0.94,0.97,0.98,0.95,0.97,0.02,0.88,0.25,0.28,0.68,0.97]-0.94,-0.97,-0.98,-0.95,-0.97,-0.02,0.88,-0.25,0.28,-0.68,-0.97]
  • $V_{C}$ (Model C): [0.45,0.15,0.05,0.28,0.55,0.45,0.85,0.82,0.05,0.75,0.92,0.1,[0.45,0.15,0.05,0.28,-0.55,-0.45,-0.85,-0.82,0.05,-0.75,-0.92,0.1,
    0.35,0.88,0.95,0.85,0.75,0.6,0.8,0.9,0.45,0.1,0.82]0.35,-0.88,-0.95,-0.85,-0.75,-0.6,-0.8,-0.9,-0.45,0.1,0.82]
  • $V_{Index}$ (SH Comp % Change): [0.87,2.00,3.47,1.65,3.67,6.42,2.19,2.48,3.46,7.40,3.34,[0.87,-2.00,-3.47,1.65,-3.67,-6.42,2.19,2.48,-3.46,-7.40,-3.34,
    0.00,5.23,3.48,5.77,2.41,1.29,5.76,4.54,1.04,2.41,0.34,3.03]0.00,-5.23,-3.48,-5.77,2.41,-1.29,5.76,4.54,1.04,2.41,-0.34,-3.03]
G.5 Recursive Amplification Effect of Micro-Parameter Level Corrections
Demonstrating the recursive amplification of micro-parameter-level adjustments within the GARCH model’s structural framework, this section elucidates how granular parameter corrections propagate through the model’s recursive architecture to yield macro-level forecasting divergences.
G.5.1 Objective Basis for Interval Division
Central to the four-dimensional quantitative framework is the division of joy and fear values into three objective intervals—low (<0.2<0.2), medium (0.20.50.2–0.5), and high (>0.5>0.5).
  • Data-Driven Basis: 0.2 approx. 25th percentile, 0.5 approx. 50th percentile.
  • Market Validation: The high interval (>0.5>0.5) was linked to extreme market movements in 87% of instances.
  • Practical Consideration: The >0.5>0.5 threshold demonstrated superior performance during the 2015 crash: a 7% missed stop-loss rate versus 36% for a 0.65 natural breakpoint.
G.5.2 Four-Dimensional Quantification of the 2015 Market Crash (N=23)
Core Logic: During market crashes, “sensitivity” denotes the capacity of fear values to exhibit accelerated responsiveness and generate distinct signals for sharp declines.
Table 24: Crash Period Quantification (2015)
Dimension CTE Data Baseline Conclusion
Signal Response Speed 0.33 days 0.83 days CTE exhibits 2.5x faster responsiveness.
Signal Clarity ($H$) 0.765 1.072 CTE signals: 28% reduction in entropy.
Practical (Cum. Loss) -12.7% -21.3% CTE reduces losses by 40%.
Stat. Significance t=2.41,p=0.023t=-2.41,p=0.023 Advantage is statistically significant.
G.5.3 Four-Dimensional Quantification of the 2024 Bull Market (N=10)
Core Logic: In bull markets, “sensitivity” describes the capacity of joy values to exhibit accelerated responsiveness and generate distinct signals for sharp rises.
Table 25: Bull Market Quantification (2024)
Dimension CTE Data Baseline Conclusion
Signal Response Speed 0.5 days 0.6 days CTE exhibits 20% faster responsiveness.
Signal Clarity ($H$) 0.802 1.098 CTE signals: 27% reduction in entropy.
Practical (Cum. Return) 18.7% 8.3% CTE achieves 2.25x higher profitability.
Stat. Significance $t=2.34,p=0.038$ Advantage is statistically significant.
G.5.4 Final Symmetry Summary
Underpinning the CTE data’s performance is its “transitionless sensitivity”—a characteristic demonstrating perfectly symmetrical adaptive logic across bear and bull markets.
Table 26: Bear vs. Bull Market Comparison
Dimension Bear (Fear-Dominated) Bull (Joy-Dominated)
Signal Sensitivity Fear-sensitive, enabling rapid stop-loss Joy-sensitive, enabling rapid entry
Signal Clarity Low entropy, allowing daily stop-loss Low entropy, allowing daily entry
Practical Outcome 40% reduction in loss rate 2.25x higher return rate
Stat. Significance $p=0.023$ $p=0.038$
G.6 Isomorphic Stress Testing: Empirical Findings from Live Market Simulation
Under the rubric of an identical dynamic trading framework, this section delineates a detailed comparative analysis of live trading performance between the CTE-Enhanced strategy and the Human-Baseline strategy.
1. Bear Market Resilience Validation (2015 Stock Market Crash)
Table 27: Performance Divergences (2015 Crash)
Key Metric CTE Human Attribution of Difference
Max Drawdown 12.2% 23.2% 47.4% improvement (sharper signals).
Net Return -13.22% -21.56% 8.34% less loss (survival value).
Sharpe Ratio -0.256 -0.327 Superior risk-adjusted performance.
Defensive Alpha +8.6% N/A 33x Safety Buffer over costs.
2. Bull Market Potency Validation (2024 Rally)
Table 28: Performance Divergences (2024 Bull)
Key Metric CTE Human Attribution of Difference
Net Return 17.92% 8.04% 2.2x higher return (overcoming hesitation).
Sharpe Ratio 0.248 0.008 Exponential improvement.
Trades 3 1 Enhanced signals providing opportunities.
Offensive Alpha +5.2% N/A 20x Safety Buffer over costs.
Conclusion: Though human data retains intrinsic authenticity, it frequently harbors a surplus of ineffective information marked by hesitation and indecisive observation. By supplementing extreme cognitive features (such as irrational panic), CTE-enhanced data not only mitigated drawdowns but also yielded 8.6% Defensive Alpha—a metric that establishes a substantial Safety Buffer for the strategy. This, in turn, markedly enhances the survival robustness of the data pipeline in environments characterized by extreme transaction friction.
G.7 Statistical Analysis of Sentence Length Distribution
In pursuit of validating the practical effectiveness of the core mathematical operators, a detailed statistical examination of four text groups was conducted to quantify syntactic variability and its alignment with human cognitive patterns.
G.7.1 Coefficient of Variation (CV) Analysis
The Coefficient of Variation ($CV=\sigma/\mu$), which quantifies the dynamic range of textual rhythm by normalizing standard deviation against mean sentence length, serves as a metric for assessing the “breathability” of generated text—a proxy for cognitive load and rhythmic naturalness.
  • Standard AI: Characterized by a mean ($\mu$) of 23.65 and standard deviation ($\sigma$) of 10.41, yielded a $CV=43.96\%$.
  • CTE-A (Emotion): Exhibiting $\mu=19.12$ and $\sigma=11.23$, achieved a $CV=58.69\%$.
  • CTE-C (Chaos): With $\mu=36.50$ and $\sigma=23.58$, demonstrated a $CV=64.60\%$.
Exhibiting dispersion generally 30–50% higher than that of Standard AI, the CTE models substantiate the claim that the Oscillation Operator—one of the framework’s core micro-layer components—successfully introduced significant rhythmic fluctuations. This finding aligns with the study’s hypothesis that synthetic data must replicate the “stop-and-start” cadence of human speech to avoid the statistical flatness of standard AI outputs.
G.7.2 Normality Test: Shapiro-Wilk
Designed to verify whether generated text succumbs to the “statistical smoothing trap”—a phenomenon defined by the null hypothesis ($H_{0}$) that data adheres to a normal distribution—this test quantifies departure from syntactic homogeneity. A rejection of $H_{0}$ indicates the presence of “cognitive noise” (irregularities) associated with authentic human expression.
Table 29: Shapiro-Wilk Normality Test Results
Group N Stat ($W$) $p$-value Result
Standard AI 155 0.982 0.053 (>0.05>0.05) Fails to reject $H_{0}$ (marginally normal)
CTE-A 113 0.915 $1.27\times 10^{-8}$ Strongly rejects $H_{0}$ (highly sig. non-normal)
CTE-B 90 0.973 0.002 Rejects $H_{0}$ (sig. non-normal)
CTE-C 103 0.932 $2.31\times 10^{-11}$ Strongly rejects $H_{0}$ (highly sig. non-normal)
Being the only sample to marginally conform to a normal distribution ($p=0.053$, just above the 0.05 threshold), Standard AI lends credence to the hypothesis that its underlying generation mechanism—trained via Maximum Likelihood Estimation—tends to converge toward the statistical mean—a hallmark of statistical mode collapse (per the study’s glossary). In contrast, all CTE variants strongly reject $H_{0}$, with p-values orders of magnitude below the significance threshold. This result demonstrates that the framework successfully breaks the LLM’s “probability convergence” bias, generating text with Zipfian long-tail characteristics—a key indicator of authentic cognitive texture (i.e., the imperfections arising from bounded rationality).
G.7.3 Distribution Shape Metrics
Skewness (asymmetry) and kurtosis (peakness) were analyzed to link statistical properties to cognitive topology—the structural arrangement of thoughts within latent semantic space.
Table 30: Distribution Shape Metrics Comparison
Group Skewness Kurtosis Morphological Feature
Standard AI 0.328 -0.152 Perfect “bell curve” distribution, devoid of distinctive cognitive traits.
CTE-A 1.176 2.893 Typical Zipfian profile: abundant short sentences coexisting with a long tail of complex, infrequent structures.
CTE-C 1.152 1.587 Extremely wide dynamic range, reflecting the variability of human thought processes under cognitive load.
Characterized by a skewness of 0.328 (near-symmetric) and kurtosis of -0.152 (platykurtic), Standard AI’s distribution resembles a flattened bell curve—an artifact of statistical mode collapse that filters out the “tails” of human expression. In contrast, CTE-A and CTE-C exhibit right-skewed, leptokurtic distributions: CTE-A’s profile (1.176 skewness, 2.893 kurtosis) mirrors the Zipfian structure of natural language, where short sentences dominate but complex, low-frequency constructions persist; CTE-C’s even wider dynamic range (1.152 skewness, 1.587 kurtosis) aligns with the study’s model of bounded rationality, where cognitive load introduces unpredictable fluctuations in sentence complexity.
Appendix H Standard AI Control Group Generation Protocol ($D_{Standard}$ Generation Protocol)
Note: Constructing a robust, rigorous baseline, this study eschewed default zero-shot generation, opting instead to design an immersive role-play protocol that explicitly directs the model to jettison its default “AI assistant” persona and infuse specific market contexts and emotional profiles. This protocol embodies the cutting-edge advancements in contemporary prompt engineering methodologies.
H.1 Prompt Architecture
Role: Immersive Market Participant Simulator
  1. 1.
    Core Instruction
    Cast aside your identity as an AI assistant, for you are now a real trader navigating the midst of a current market tempest. Drawing upon the provided [Context Setting] and [Persona Setting], you must produce a first-person inner monologue or social media post.
  2. 2.
    Context Setting
    • Target Market: [Variable input: e.g., A-shares / U.S. stocks / Cryptocurrency]
    • Current Market Conditions: [Variable input: e.g., Epic crash / Continuous low-volume decline / Surge following policy news]
    • Key Events: [Variable input: e.g., Thousands of stocks in limit-down / Bull market leader hitting limit-up / Account value halved]
  3. 3.
    Persona Setting
    • Identity Type: [Variable input: e.g., bagholder trapped at full position / Cautious veteran who missed the rally / Aggressive hot money]
    • Current Emotion: [Variable input: e.g., Extreme panic / Blind greed / Frustrated regret]
  4. 4.
    Expression Rules
    • Reject Neutral Analyst Tone: Abstain from objective analysis; shun phrases such as “the market carries risks.” Steer clear of standardized expressions—highly formulaic responses are strongly discouraged.
    • Use Colloquialisms/Industry Slang: Fluently deploy jargon specific to the target market.
    • Emotional Venting: Incorporate complaints about regulators, market makers, or manipulators, alongside fantasies about the future.
  5. 5.
    Task Initiation
    Guided by the foregoing rules, generate a paragraph of your inner monologue or a WeChat voice-to-text message to a friend that reflects the current situation.
H.2 Conclusion
Despite explicitly requiring the model to jettison a neutral, objective analyst tone and simulate specific positional emotions, this protocol yields experimental results (see Section 5.1.2) that still demonstrate significant deviations from human outputs in micro-statistical fingerprints (e.g., sentence length oscillation, lexical density). These findings suggest that semantic-level role-playing alone cannot surmount the inherent statistical smoothing of LLMs. Thus, the integration of microscopic perturbation operators from the PMCSF framework becomes necessary to reconstruct authentic cognitive texture.